### v8

# Digital Durable Goods Monopoly

Abstract
Abstract
This paper develops a model of dynamic monopoly for digital durable goods characterized by two features: free disposability on the buyer side and zero marginal cost on the seller side. We show that when parties are sufficiently patient, the set of equilibrium seller payoffs contains the interval from the lowest-type buyer’s efficient surplus to the static monopoly commitment payoff, subject to the constraint that the lowest-type buyer receives an efficient allocation with probability one. The reason is that when allocative efficiency is set-valued rather than single-valued, Coasian forces do not generally pin down continuation values and therefore do not select a unique payoff outcome. These results show how structural properties of digital goods generate market indeterminacy.
ntroduction
The high profitability of business-to-business digital goods markets makes them an important object of study [^1] See evidence in Appendix C.1. Business-to-business (B2B) digital goods, such as enterprise software, curated datasets, and access to APIs or AI models, are rarely purchased as standardized off-the-shelf products. Instead, they are acquired through negotiated contracts that bundle multiple attributes, including capacity, features, tiers, service levels, and price, into a single offer that the buyer either accepts or rejects. The bargaining process typically concerns custom pricing and licensing configurations, in which the vendor maintains a unified codebase while offering distinct modules or tiers of functionality. For example, a client in a regulated industry may require an “Advanced Compliance” module, whereas another client may view such features as unnecessary.
Two contractual features are salient in many digital B2B markets. First, buyers can underutilize purchased capacity or functionality and leave unused portions idle at essentially no cost, and realized usage is often non-contractible (Corrao et al., 2023). For example, a firm purchasing a dataset from a data broker such as Acxiom may query only the subset of consumer profiles relevant to a campaign. A corporation licensing enterprise software from Salesforce may leave entire modules (e.g., “Salesforce Maps”) idle. Second, once a digital product has been developed, expanding contractual access, such as enabling modules, granting seats, or raising usage limits, is close to costless at the margin relative to the fixed development cost (Quah, 2003; Goldfarb and Tucker, 2019). For data brokers such as Acxiom or Experian, after consumer records are aggregated and refined, providing filtered subsets or curated segments to clients involves minimal cost. Likewise, after firms such as Microsoft or OpenAI develop their core platforms, extending access to different feature sets, tiers, or API limits requires no significant resource expenditure. These features induce a distinct strategic structure. The buyer can accept an allocation that exceeds the efficient utilization level and discard the excess, while the seller can increase contracted access without incurring incremental production costs along the contractual allocation dimension. As a result, allocative efficiency is non-unique in a way that does not arise in standard physical-goods environments.
These negotiations also lie at the center of a classic theoretical problem: a monopolist with limited commitment selling a durable good to a buyer with private information. In canonical durable-goods monopoly models, this lack of commitment disciplines market power and can drive outcomes toward efficiency in the patient limit, as formalized by the Coase conjecture (Coase, 1972; Fudenberg et al., 1985; Gul et al., 1986). This paper uses a digitally motivated contracting environment to identify a boundary of this logic. When allocative efficiency is set-valued rather than single-valued, Coasian forces continue to impose constraints, but they do not generally pin down continuation values and hence do not select a unique payoff outcome.
Throughout the paper, “efficiency” refers to allocative efficiency in the allocation or quality dimension, conditional on trade, rather than efficiency in the timing dimension. [^2] With quality choice, equilibrium delay patterns are difficult to characterize in general (See an indeterminacy result in Appendix B.4). Accordingly, the main results focus on the set of seller payoffs sustainable in equilibrium and how this set changes when allocative efficiency is non-unique. The central question is therefore the following. What constraints does Coasian discipline impose when the set of efficient allocations is large, and which payoff outcomes can be sustained?
We formalize these questions in a dynamic monopoly model with one-sided private information. The buyer has a privately known type that captures the marginal benefit of utilization, or equivalently the negative of a marginal implementation cost. In each period, the seller makes a single take-it-or-leave-it offer consisting of an allocation–price pair, and the buyer either accepts or rejects. The allocation is interpreted as a contractual allocation, such as access rights, enabled features, seats, capacity, or precision, while realized usage is chosen by the buyer after purchase. Formally, the buyer may discard unused portions at zero cost (free disposability). On the seller side, we assume zero marginal cost along the contractual allocation dimension, so the seller’s profit equals the transfer. The single-offer protocol isolates intertemporal screening. It provides a reduced form for many B2B private-offer negotiations, in which the contracting object is a bundled proposal rather than a posted menu; Appendix C.2 discusses institutional support. Allowing within-period menus would reintroduce intratemporal screening and reduce the dynamic problem to a static menu benchmark, which we analyze separately in Appendix B.3.
The main result is a Folk-Theorem-type attainability statement for seller payoffs. Let $u_{\text{min}}$ denote the surplus of the lowest-type buyer under the efficient allocation. Let $\pi^{e}(F)$ denote the seller’s optimal static payoff subject to the additional constraint that the lowest-type buyer receives an allocatively efficient allocation with probability one. We show that for every $\epsilon>0$, there exists $\underline{\delta}$ such that for all $\delta\geq\underline{\delta}$, the set of equilibrium seller payoffs contains the interval
$u_{\text{min}},\ \pi^{e}(F)-\epsilon].$
Moreover, these high-payoff equilibria can be constructed to satisfy an on-path Markov restriction on buyer behavior, a refinement that often collapses multiplicity in canonical Coasian settings. In particular, we do not claim that $\pi^{e}(F)$ is a universal upper bound across all equilibria in the general quality-choice environment. Appendix B.1 provides upper-bound results for broad classes of equilibria.
Free disposability and zero marginal contractual cost imply a sharp structural property. Each type has a unique efficient utilization level, but any contractual allocation above is allocatively efficient for that type because the buyer can discard slack. Consequently, the seller can change the contracted allocation path without reducing allocative efficiency. This decouples two roles that coincide in canonical Coasian models: (i) delivering efficiency and (ii) pinning down continuation values. When efficiency is essentially single-valued, it tightly restricts feasible continuation paths and often yields sharp predictions. When allocative efficiency is set-valued, the seller can select allocation paths that generate different continuation values while respecting the incentive constraints on the marginal type. This flexibility sustains the interval of attainable payoffs.
The role of “zero marginal cost” (together with free disposability) should be understood as a structural boundary rather than an innocuous approximation. The payoff indeterminacy disappears when allocative efficiency becomes single-valued, e.g., without free disposability or with strictly increasing marginal costs. Section 5 provides detailed results in discrete-type environments.
Thus, in digital contracting environments, the negotiated outcomes may be more sensitive to norms and bargaining practices, and less tightly pinned down by fundamentals than in canonical models.
Our set-valued efficiency perspective also helps clarify why Ausubel and Deneckere (1989b) obtain a folk theorem result, as a case in which the lowest type’s efficient set is non-singleton. A detailed discussion is deferred to Section 4.4.
Roadmap.
Section 2 presents the model. Section 3 characterizes static benchmarks. Section 4 establishes the Coasian equilibrium and the Folk-Theorem-type attainability result. Section 5 studies cases with unique efficient allocations. Section 6 discusses applications. A detailed literature review is deferred to Section 7. Section 8 concludes. The appendices provide proofs, additional results, and institutional discussion.
odel
2.1 Primitives
The buyer is characterized by a privately known type $\theta\in\Theta=[\underline{\theta},\overline{\theta}]$, drawn from a cumulative distribution function $F(\cdot)$. [^3] In the baseline model, buyer types are drawn from a continuum to simplify the static commitment benchmark. We analyze finite-type distributions in Section 5 and Appendix B. The type $\theta$ may take negative values; for instance, it can be interpreted as c-c, representing marginal implementation or utilization costs.
A single good is available for trade. The allocation space is $X=[\underline{x},\overline{x}]$, where $\underline{x}>0$ denotes the minimal technologically feasible allocation. The variable $x$ may represent either the quantity or the quality of the good. We interpret $x$ as a contractual allocation (access/rights/capacity), while realized usage $x^{\prime}$ is chosen by the buyer and is not directly contractible.
Given an allocation–price pair $(x,p)$, the buyer has quasilinear preferences over consumption and transfers. The utility is
$u(x,\theta)-p,$
where the gross utility $u(x,\theta)$ incorporates the free disposability feature:
$u(x,\theta)=\max_{0\leq x^{\prime}\leq x}\left[v(x^{\prime})+\theta x^{\prime}\right].$
This formulation implies that although the buyer purchases an allocation $x$, she optimally selects an actual consumption level $x^{\prime}\in[0,x]$. Any excess allocation $x-x^{\prime}$ is discarded without cost. Under zero marginal cost along the contractual allocation dimension, the seller’s profit equals $p$.
Observe that for any fixed allocation $x\in X$, the utility function $u(x,\cdot)$ is non-decreasing in the buyer’s type $\theta$.
We maintain the following assumptions throughout the analysis.
Assumption 1.
The function $v(\cdot)$ is continuously differentiable and strictly concave on $0,\overline{x}]$, with $v(0)=0$.
Assumption 1 captures strictly diminishing marginal returns to consumption. To establish the efficiency benchmark, we characterize the utility-maximizing consumption level for each buyer type.
Lemma 1 (Unique Efficient Consumption).
For each type $\theta$, the unique consumption level that maximizes utility, denoted $x^{e}(\theta)$, is
$x^{e}(\theta)=\begin{cases}v^{\prime-1}(-\theta),&\text{if }v^{\prime-1}(-\theta)\in[0,\overline{x}],\\[6.0pt] 0,&\text{if }v^{\prime}(0)+\theta<0,\\[6.0pt] \overline{x},&\text{if }v^{\prime}(\overline{x})+\theta>0.\end{cases}$
Furthermore, $x^{e}(\theta)$ is continuous and non-decreasing in $\theta$.
Lemma 2 (Efficiency of Allocations).
An allocation $x$ is efficient for type $\theta$ if and only if $x\geq x^{e}(\theta)$.
We omit the proofs of Lemma 1 and Lemma 2, as they follow from the strict concavity of $v(\cdot)$ and the free disposability property.
We next impose a standard regularity condition on the distribution of buyer types.
Assumption 2.
Buyer types are distributed according to a cumulative distribution function $F(\cdot)$ with a density $f(\cdot)$ that is strictly positive and bounded (i.e., $0<m\leq f(\theta)\leq M$ for all $\theta\in\Theta$). The virtual valuation function
$J(\theta)\equiv\theta-\frac{1-F(\theta)}{f(\theta)}$
is strictly increasing.
Assumption 2 corresponds to the monotone hazard rate condition, which ensures regularity of the screening problem.
Finally, we introduce a technical assumption that simplifies the exposition.
Assumption 3 (No-Exclusion at the Efficient Threshold).
$v(x^{e}(\theta))+\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right)x^{e}(\theta)\geq 0\quad\text{for all }\theta\in\Theta.$
Assumption 3 requires that the virtual surplus at the efficient consumption level be nonnegative for every type $\theta$. This assumption is imposed for equilibrium existence in the continuum‑type, continuous‑action baseline and does not drive the results; Appendix B shows the main constructions in a discretized action space without Assumption 3.
In particular, since the gross surplus $v(x^{e}(\theta))+\theta x^{e}(\theta)$ is non-decreasing in $\theta$, we have
$v(x^{e}(\theta))+\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right)x^{e}(\theta)$ $\geq v(x^{e}(\underline{\theta}))+\underline{\theta}x^{e}(\underline{\theta})-\frac{1-F(\theta)}{f(\theta)}x^{e}(\theta)$
$\geq v(x^{e}(\underline{\theta}))+\underline{\theta}x^{e}(\underline{\theta})-\frac{1}{m}\overline{x}.$
Thus, a sufficient condition for Assumption 3 to hold is
$v(x^{e}(\underline{\theta}))+\underline{\theta}x^{e}(\underline{\theta})\;\geq\;\frac{\overline{x}}{m}.$
2.2 Timing and Solution Concept
Time is discrete and indexed by $t=0,1,\dots$, with a common discount factor $\delta\in[0,1)$. In each period $t$, conditional on trade not having occurred, the seller offers an allocation–price pair $(x_{t},p_{t})\in[\underline{x},\overline{x}]\times\mathbb{R}$.
Upon observing $(x_{t},p_{t})$, the buyer chooses to accept ($a_{t}=1$) or reject ($a_{t}=0$). A seller history at the beginning of period $t$ is
$h_{t}:=(x_{i},p_{i},a_{i})_{i=0}^{t-1},$
with $H_{t}$ the set of such histories and $H:=\cup_{t=0}^{\infty}H_{t}$ the set of all seller histories. The buyer’s history $\hat{h}_{t}$ consists of $h_{t}$ and the current offer $(x_{t},p_{t})$. Let $\hat{H}$ denote the set of all buyer histories.
Acceptance terminates the game [^4] This is consistent with a durable access right (or a negotiated contract whose terms remain fixed over the relevant horizon).. If the buyer accepts in period $t$ ($a_{t}=1$), the buyer’s payoff is
$\delta^{t}\bigl[u(x_{t},\theta)-p_{t}\bigr],$
and the seller’s payoff is
$\delta^{t}p_{t}.$
A behavioral pure strategy for the buyer is a measurable function
$\alpha:\hat{H}\times[\underline{\theta},\overline{\theta}]\to\{0,1\},$
[…]
The seller’s behavioral strategy is a measurable function
$\sigma:H\to\Delta\bigl([\underline{x},\overline{x}]\times\mathbb{R}\bigr),$
where $\Delta(\cdot)$ denotes the set of probability distributions over the outcome space.
A Perfect Bayesian Equilibrium (PBE) consists of a strategy profile $(\sigma,\alpha)$ and a system of beliefs over the distribution of active buyer types, satisfying:
  1. (i)
    Sequential Rationality: Given the beliefs, the strategies maximize expected payoffs at every history.
  2. (ii)
    Consistency: Beliefs are updated using Bayes’ rule whenever possible.
tatic Benchmark
We analyze a static benchmark to establish an upper bound on the seller’s payoff in the dynamic setting. In this benchmark, the seller commits to a direct mechanism that assigns each buyer type a probability of trade, a lottery over allocations, and a transfer. Formally, a mechanism $\mathcal{M}$ is a mapping
$\mathcal{M}:[\underline{\theta},\overline{\theta}]\to[0,1]\times\Delta([\underline{x},\overline{x}])\times\mathbb{R},$
where $\Delta([\underline{x},\overline{x}])$ is the set of probability distributions over the allocation space $\underline{x},\overline{x}]$. For a type $\theta$, we denote the mechanism components by $(q(\theta),\tilde{x}(\theta),p(\theta))$, where $q(\theta)$ is the probability of trade, $\tilde{x}(\theta)$ is the random allocation, and $p(\theta)$ is the expected transfer.
A mechanism satisfies Incentive Compatibility (IC) if every buyer type $\theta$ weakly prefers its assigned allocation to that of any other type $\theta^{\prime}$:
$q(\theta)\,\mathbb{E}\bigl[u(\tilde{x}(\theta),\theta)\bigr]-p(\theta)\ \geq\ q(\theta^{\prime})\,\mathbb{E}\bigl[u(\tilde{x}(\theta^{\prime}),\theta)\bigr]-p(\theta^{\prime}),\quad\forall\theta,\theta^{\prime}\in\Theta.$
It satisfies Individual Rationality (IR) if each type weakly prefers participation to the outside option:
$q(\theta)\,\mathbb{E}\bigl[u(\tilde{x}(\theta),\theta)\bigr]-p(\theta)\ \geq\ 0,\quad\forall\theta\in\Theta.$
The seller’s optimization problem is therefore
$\max_{\mathcal{M}}\ \int_{\underline{\theta}}^{\overline{\theta}}p(\theta)f(\theta)\,\mathrm{d}\theta,$
subject to the IC and IR constraints.
This static benchmark is similar to the monopoly quality differentiation problem analyzed by Mussa and Rosen (1978). Although we admit lotteries over allocations due to the nonlinearity of $u(\cdot,\theta)$, [^6] Given the strict concavity of $v(\cdot)$, deterministic allocations strictly dominate lotteries in the optimal mechanism. standard mechanism design techniques yield the following characterization.
Lemma 3 (Maximum Static Profit).
The revenue-maximizing mechanism assigns each buyer type $\theta$ the allocation
$x^{*}(\theta)\in\operatorname*{arg\,max}_{x\in[\underline{x},\overline{x}]}\left\{v(x)+x\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right)\right\}.$
Each type $\theta$ consumes
$\min\{x^{e}(\theta),\,x^{*}(\theta)\},$
and pays
$p(\theta)=u(x^{*}(\theta),\theta)-\int_{\underline{\theta}}^{\theta}\frac{\partial u}{\partial z}(x^{*}(z),z)\,\mathrm{d}z.$
The seller’s optimal payoff is denoted by $\pi(F)$.
The profit $\pi(F)$ from the static framework provides an upper bound on the seller’s revenue in the dynamic environment. This follows from the “revelation principle” for dynamic games (e.g., Ausubel and Deneckere, 1989a): any outcome supported by a seller’s strategy and a buyer’s best response in the dynamic game induces a distribution over allocations and transfers that can be replicated by an incentive-compatible direct mechanism. Thus, even with full commitment, the seller cannot exceed the payoff of the optimal static mechanism.
Corollary 1.
No seller strategy, coupled with a buyer best response, can yield a payoff strictly exceeding $\pi(F)$.
We now examine a constrained static benchmark that forms the basis for the Folk Theorem. In this benchmark, the lowest buyer type $\underline{\theta}$ must receive an efficient allocation with probability one. Formally, we consider the set of Incentive Compatible (IC) and Individually Rational (IR) mechanisms satisfying
[…]
This constrained benchmark matches the efficiency requirement that emerges in our equilibrium constructions for the lowest type. Adapting the analysis from Lemma 3 yields the following characterization.
Lemma 4 (Maximum Profit with Lowest-Type Efficiency).
The revenue-maximizing mechanism subject to these constraints assigns each type $\theta$ the allocation
$x^{*}(\theta)\in\operatorname*{arg\,max}_{x\in[\max\{x^{e}(\underline{\theta}),\underline{x}\},\overline{x}]}\left\{v(x)+x\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right)\right\}.$
Each type $\theta$ consumes
$\min\{x^{e}(\theta),\,x^{*}(\theta)\},$
and the associated transfer is
$p(\theta)=u(x^{*}(\theta),\theta)-\int_{\underline{\theta}}^{\theta}\frac{\partial u}{\partial z}(x^{*}(z),z)\,\mathrm{d}z.$
We denote the seller’s optimal payoff under these constraints by $\pi^{e}(F)$.
Typically, the seller distorts allocations downward to maximize virtual surplus. The efficiency requirement for the lowest type forces the seller to assign at least $x^{e}(\underline{\theta})$ to type $\underline{\theta}$, which, by IC, establishes a lower bound of $x^{e}(\underline{\theta})$ for the entire type space. In the unconstrained benchmark (Lemma 3), the optimal mechanism would involve distortions below this bound whenever the technological constraint $\underline{x}$ permits.
The following corollary is immediate.
Corollary 2.
The inequality $\pi^{e}(F)\leq\pi(F)$ holds by definition. Moreover, $\pi^{e}(F)=\pi(F)$ if and only if the technological minimum satisfies $\underline{x}\geq x^{e}(\underline{\theta})$.
ain Results
This section presents the main results. We begin by characterizing a Coasian equilibrium, which yields a lower bound on the seller’s payoff. In this equilibrium, the lack of commitment drives the seller’s profit to the lowest type buyer’s efficient surplus.
4.1 The Coasian Equilibrium
We construct a skimming equilibrium in which every offer—on-path and off-path—is accepted by an upper interval of buyer types. Define
$u_{\min}:=v(x^{e}(\underline{\theta}))+\underline{\theta}x^{e}(\underline{\theta})$
as the surplus of the lowest-type buyer under the efficient allocation.
Theorem 1 (Coasian Equilibrium).
For any discount factor $0\leq\delta<1$, there exists an equilibrium in which the seller clears the market immediately at $t=0$ by offering $(\overline{x},u_{\min})$, and obtains revenue equal to $u_{\min}$.
The logic underlying this result is that the seller can implement efficient allocations for all buyer types within a single period [^7] The importance of this property will become clear in Section 5, where we examine the case without free disposability. The payoff as low as $u_{\text{min}}$ cannot, in general, be sustained in equilibrium without free disposability. . Although efficient consumption levels are type-specific, the seller can satisfy efficiency simultaneously by offering the maximal allocation $\overline{x}$. With free disposal, each buyer type attains its efficient consumption from this offer. On the equilibrium path, because every buyer consumes their efficient level, Assumption 3 implies that immediate market clearing is optimal within the constructed strategy profile.
Should the seller deviate by offering a distinct allocation, buyers anticipate that the “quality” component of the offer will persist (i.e., no future downward adjustments). Under these anticipations of future strategies, price becomes the sole screening instrument. Consequently, any deviation from the efficient allocation reduces total surplus and therefore lowers the seller’s profit.
These strategies are sustainable in equilibrium because, on the constructed path, the seller clears the residual market in the subsequent period regardless of the initial deviation. At the point of market clearing, the seller is indifferent across all allocations in $x^{e}(\underline{\theta}),\overline{x}]$, as revenue is determined by the surplus of the lowest type $\underline{\theta}$. Thus, immediate market clearing at $t=0$ maximizes the seller’s payoff within the constructed strategy profile.
This outcome mirrors the standard Coasian benchmark in the following sense: every buyer type receives an efficient allocation with no delay, and the seller’s profit is driven to the surplus of the lowest buyer type. Since $u_{\min}<\pi(F)$, this payoff lies strictly below the static commitment benchmark.
4.2 Departure from the Single-Quality Benchmark
Before presenting the Folk Theorem, we describe how this setting diverges from the Coasian bargaining framework with a single quality level. When the seller offers allocation–price pairs $(x_{t},p_{t})$ rather than scalar prices $p_{t}$, the skimming property need not hold. In particular, one cannot assume a priori that the seller’s posterior beliefs are upper truncations of the prior distribution—that is, that the set of active buyers always forms a connected interval $\underline{\theta},\theta]$. As a result, the highest remaining type is generally insufficient as a state variable for dynamic programming.
This structural difference invalidates standard backward-induction arguments. In the canonical Coasian setting, backward induction relies on the existence of a uniform upper bound $T(\delta)$ on the time to market clearing, which in turn depends on the geometric property that the posterior support remains an interval. In the present context, however, the set of remaining types may fragment into an arbitrary finite union of disjoint intervals. Consequently, standard recursive techniques are infeasible. In addition, monotone comparative statics (Milgrom and Shannon, 1994) are not guaranteed even if the standard skimming property holds, due to the presence of an additional allocation dimension. Therefore, the proof of Theorem 1 adopts a constructive approach: we postulate candidate equilibrium strategies and verify that they constitute an equilibrium.
4.3 The Folk Theorem
Using the Coasian equilibrium characterized in Theorem 1, we establish the Folk Theorem.
Theorem 2 (The Folk Theorem).
For every $\epsilon>0$, there exists a discount factor $\underline{\delta}$ such that for all $\delta\geq\underline{\delta}$,
$u_{\min},\ \pi^{e}(F)-\epsilon]\ \subseteq\ SE(\delta),$
where $SE(\delta)$ is the set of seller payoffs sustainable in equilibrium.
If $\underline{x}\geq x^{e}(\underline{\theta})$, then $\pi^{e}(F)=\pi(F)$ by Corollary 2. We therefore obtain the following corollary.
Corollary 3.
Suppose $\underline{x}\geq x^{e}(\underline{\theta})$. For any discount factor $\delta$,
$u_{\min},\ \pi^{e}(F)]\ \supseteq\ SE(\delta).$
We discuss whether $\pi^{e}(F)$ constitutes the upper bound on all equilibrium payoffs in more general settings in Appendix B.
The high-payoff equilibrium is constructed as follows. In each period $t$, the seller offers the allocation $x_{t}=x^{*}(\theta_{t})$ (as defined in Lemma 4) associated with a cutoff type $\theta_{t}$. The price $p_{t}$ is chosen so that type $\theta_{t}$ is indifferent between accepting immediately and waiting one period. All types $\theta\geq\theta_{t}$ accept. As $\delta\to 1$ and the step size $\theta_{t}-\theta_{t+1}\to 0$, this sequence of offers converges to the menu characterized in Lemma 4.
If the seller deviates from this sequence, play reverts immediately and permanently to the Coasian equilibrium in Theorem 1. This threat is credible because the Coasian outcome is itself an equilibrium. The construction is analogous to the reputational equilibria in Ausubel and Deneckere (1989b), but our analysis explicitly addresses the gap case and incorporates finite-time market clearing. In the gap case without quality differentiation, the Coase conjecture obtains, while in Ausubel and Deneckere (1989b) the market does not clear in finite time.
The high-payoff equilibrium can be interpreted through an endogenous reputation perspective. Initially, the buyer believes that the seller is a strong type who will adhere to the equilibrium price path. If a deviation from this path occurs, the buyer infers that the seller is weak and will henceforth follow the equilibrium characterized by Theorem 1. The resulting loss of reputation deters the seller from deviating.
4.4 Comparison
In this section, we compare Theorem 2 to the benchmark setting without quality differentiation. We show that our Folk Theorem is both analogous to and distinct from the established benchmark.
Relation to the Established Folk Theorem.
In the framework of Ausubel and Deneckere (1989b), where the lowest buyer valuation is normalized to zero (the “no-gap” case), the equilibrium payoff set extends to the full commitment payoff. Our analysis provides a complementary perspective on this result. In the setting of Ausubel and Deneckere (1989b), the only allocation dimension is the discounted probability of trade. When the lowest type has valuation zero in the no-gap case, the efficient timing of trade is set-valued for this type, including infinite delay. Thus, under limited commitment, once the lowest type is allocated infinite delay, the allocations for other types may be arbitrary, provided they are monotone in type. In this context, the upper bound of the Folk Theorem can be interpreted as the maximum profit achievable subject to the constraint that the lowest-type buyer receives an efficient time of trade. In their setting, this constrained maximum coincides with the unconstrained static commitment payoff.
Robustness to Markov Refinements
We show that the Folk Theorem is robust to a class of Markov refinements that typically eliminate reputational equilibria and restore the Coase conjecture in standard settings. This distinguishes our environment from those benchmarks. We begin by recalling the definition of weak-Markov equilibria.
Definition 1 (Weak-Markov Property).
An equilibrium is weak-Markov if and only if the buyer’s strategy $\alpha(\cdot,\theta)$ depends solely on the seller’s current offer $(x_{t},p_{t})$.
Equivalently, a weak-Markov equilibrium is a Perfect Bayesian Equilibrium in which, for any two histories $(h,(x,p))$ and $(h^{\prime},(x,p))\in\hat{H}$,
$\alpha((h,(x,p)),\theta)=\alpha((h^{\prime},(x,p)),\theta)$
for all $\theta\in A(h,(x,p))\cap A(h^{\prime},(x,p))$, where A(h^)A($\hat${h}) denotes the set of active buyer types at history $\hat{h}$.
We now introduce a relaxation of this concept.
Definition 2 (On-Path Markov Property).
A history $h_{t}$ is on-path if it occurs with positive probability under the equilibrium strategies $(\sigma,\alpha)$ for some buyer type. An equilibrium is on-path Markov if, for any two on-path histories $h$ and $h^{\prime}$ and any offer $(x,p)$ that occurs with positive probability along the equilibrium path,
$\alpha((h,(x,p)),\theta)=\alpha((h^{\prime},(x,p)),\theta)$
for all $\theta\in A(h,(x,p))\cap A(h^{\prime},(x,p))$.
Definition 2 requires that if the seller deviates to an offer that appears later on the equilibrium path—thereby preempting a scheduled offer—buyers respond exactly as they would on the equilibrium path. Any weak-Markov equilibrium necessarily satisfies the on-path Markov property.
We establish the following corollary.
Corollary 4.
The conclusion of Theorem 2 remains valid when restricted to on-path Markov equilibria.
This result distinguishes our setting from the existing literature, in which the Coase conjecture typically persists under on-path Markov refinements. Although prior results are often stated for weak-Markov equilibria, the logic extends to the on-path Markov class. For example, in the gap case analyzed by Fudenberg et al. (1985) and Gul et al. (1986), the unique equilibrium satisfies the weak-Markov condition and therefore the on-path Markov property. In the no-gap case, Ausubel and Deneckere (1989b) show that the Folk Theorem collapses to the uniform Coase conjecture under weak-Markov (and on-path Markov) equilibria. Likewise, Nava and Schiraldi (2019) show that when the seller posts simultaneous prices for multiple goods, the Coase conjecture holds under weak-Markov—and hence on-path Markov—equilibria.
A common structural feature of these environments is that allocations are fixed, leaving price as the only screening instrument. In such settings, the on-path Markov refinement eliminates the seller’s ability to delay because the incentive to preempt future sales remains strong. In contrast, our framework allows the seller to choose both price and allocation, and the equilibrium path features decreasing allocations. This additional margin alters the incentives: accelerating sales requires offering the lower future allocation immediately, which reduces total surplus and therefore the seller’s profit. Hence, when agents are sufficiently patient, the incentive to maintain the equilibrium allocation sequence outweighs any gain from acceleration.
hen efficient allocations are unique.
In this section, we analyze the case in which efficient allocations are unique. We adopt a discrete-type framework to show that both free disposability and zero marginal cost are necessary for the payoff multiplicity highlighted by the Folk Theorem. As discussed above, the skimming property generally fails in the presence of quality differentiation, which makes a direct characterization of equilibrium strategies difficult. Addressing the problem in full generality is therefore beyond the scope of this paper.
Consider a setting with $n$ discrete types, ordered as $\theta_{1}>\theta_{2}>\dots>\theta_{n}$, each with prior probability $q_{i}>0$ such that $\sum_{i=1}^{n}q_{i}=1$.
In the baseline model with free disposability, the buyer’s utility is
$u(x,\theta)=\max_{0\leq x^{\prime}\leq x}\left[v(x^{\prime})+\theta x^{\prime}\right].$
If free disposability is absent, the buyer is constrained to consume the entire allocated amount. In this case, the utility function is
$w(x,\theta)=v(x)+\theta x.$
We first show that the market indeterminacy, that is, equilibrium multiplicity, in Theorem 2 relies on free disposability. In particular, when free disposability is absent, the equilibrium set collapses to a single point.
Proposition 1.
Suppose Assumptions 1 and 2 hold. Consider an environment with two discrete types, ordered $\theta_{1}>\theta_{2}$, each with prior probability $q_{i}>0$ such that $\sum_{i=1}^{2}q_{i}=1$. The buyer’s utility is given by
$w(x,\theta)=v(x)+\theta x.$
There exists an essentially unique equilibrium and a unique seller profit for any $\delta\in[0,1)$. Furthermore, as $\delta\to 1$, the outcome converges to efficiency, namely each type $\theta$ receives the efficient allocation $x^{e}(\theta)$ and there is no delay.
In Proposition 1, the unique equilibrium payoff converges to $\pi^{e}(F)$ in the patient limit. This occurs because, in the two-type case, the efficient outcome coincides with the static monopoly outcome subject to the constraint that the low type receives the efficient allocation. In this setting, the low type $\theta_{2}$ receives the efficient allocation. Moreover, since $\theta_{1}$ is the high type, its allocation is not subject to downward distortion in static monopoly outcome, and it therefore also receives the efficient level. On the other hand, when free disposability is present, the equilibrium payoff set spans $u_{\text{min}},\pi^{e}(F)]$. Thus, we have shown that equilibrium multiplicity relies on free disposability. Furthermore, attaining $\pi^{e}(F)$ is a special case driven by the two-type structure. In Appendix B.5, we show that $\pi^{e}(F)$ cannot be achieved in general without free disposability.
5.1 Positive Marginal Cost
Consider a setting in which the good remains freely disposable, but producing or distributing an allocation $x$ incurs a cost $c(x)$ to the seller, with $c^{\prime}(x)>0$ and $c^{\prime\prime}(x)\geq 0$ for all $x\in[\underline{x},\overline{x}]$. [^8] Consistent with the standard literature, we assume strictly increasing marginal production costs.
The seller’s profit from an allocation–price pair $(x,p)$ is given by pc(x)p-c(x). To ensure that trade with the lowest type is viable, we assume
$v(x^{e}(\underline{\theta}))+\underline{\theta}x^{e}(\underline{\theta})-c(x^{e}(\underline{\theta}))>0.$
The efficient allocation for type $\theta$ solves
$\max_{x\in[\underline{x},\overline{x}]}\left\{\max_{0\leq x^{\prime}\leq x}\big[v(x^{\prime})+\theta x^{\prime}\big]-c(x)\right\}.$
Because marginal costs are strictly increasing, $c^{\prime}(x)>0$, it is strictly suboptimal to allocate a quantity $x$ that exceeds the consumption level $x^{\prime}$. Efficiency therefore requires $x=x^{\prime}$, which reduces the problem to
$\max_{x\in[\underline{x},\overline{x}]}\big[v(x)+\theta x-c(x)\big].$
Given the convexity of the cost function $c(\cdot)$ and the strict concavity of the utility function $v(\cdot)$, the objective function is strictly concave. Consequently, the efficient allocation is unique for each buyer type.
The formal analysis is analogous to that of Propositions 1 and is therefore omitted.
Taken together, these results show that the Folk Theorem does not arise solely from quality differentiation. In this sense, they suggest that the particular indeterminacy we study is most naturally associated with digital-goods environments.
pplications
Our primary applications include digital solutions and markets for data.
Digital Solutions.
Consider the market for digital solutions, including Software-as-a-Service (SaaS) platforms (e.g., Salesforce, Microsoft), specialized APIs (e.g., Stripe, Twilio), and AI models (e.g., OpenAI). After development, these firms negotiate large-scale subscriptions with enterprise buyers. Here, $x$ represents the volume of user licenses or the set of enabled features. A buyer of type $\theta$ selects a utilization level $x^{\prime}\leq x$ to enhance productivity or support collaboration, incurring an internal implementation cost $\theta x^{\prime}$ related to training, integration, and administration. [^9] Integration and administrative costs often scale approximately linearly with utilization. The buyer’s utility is
$u(x,\theta)=\max_{0\leq x^{\prime}\leq x}\left[v(x^{\prime})+\theta x^{\prime}\right],$
where $v(\cdot)$ is strictly concave, capturing diminishing marginal returns. [^10] For example, in a CRM platform, core features provide high marginal value, while incremental tools yield lower marginal benefits.
In this setting, free disposability implies that unused licenses or features may remain idle without penalty. Zero marginal cost reflects that, once the software infrastructure is in place, providing additional licenses or feature access entails negligible marginal expense for the seller.
Consumer Data.
A second application is the market for consumer data, in which brokers such as Acxiom and Experian sell curated datasets to business clients. Consider a continuum of consumer profiles indexed by $z$ and ordered by strictly decreasing valuation, where each profile $z$ generates a gross value $v^{\prime}(z)$. [^11] Equivalently, $v^{\prime}(z)$ may be interpreted as a normalized valuation. An allocation $x$ corresponds to a marketing list containing all profiles with valuations above the threshold $v^{\prime}(x)$. Thus, the list includes profiles from the highest-value $v^{\prime}(0)$ down to the marginal profile $v^{\prime}(x)$. For a downstream firm facing a constant marginal utilization cost $c=-\theta$, the realized profit is
$u(x,\theta)=\int_{0}^{x}\left(v^{\prime}(z)-c\right)^{+}\,\mathrm{d}z.$
In this framework, datasets are organized as threshold lists, which makes $x$ a one-dimensional quality measure. This reflects the practical constraint that arbitrary sampling or recombination of individual profiles is often costly or infeasible for the seller.
Data-Driven Decision Making.
Consider a decision-maker (e.g., an investor) facing a binary choice under uncertainty regarding a state $\omega\in\{-1,1\}$, with a uniform prior. A buyer of type $\theta$ obtains payoff $\theta a\omega$ from choosing an action $a\in\{-1,1\}$ when the action matches the true state. In the absence of data, expected utility is normalized to zero. The seller offers datasets indexed by $x\in[0,1/2]$, representing the maximal available signal precision.
Possession of a dataset with precision $x$ allows the buyer to conduct a symmetric experiment that updates the prior to posteriors (1/2x, 1/2+x)($^{\prime},\,1/2+x^{\prime}$) or (1/2+x, 1/2x)($^{\prime},\,1/2-x^{\prime}$), where the realized precision $x^{\prime}\leq x$ is chosen by the buyer. Generating this signal entails a cost $C(x^{\prime})$, where $C(\cdot)$ is strictly convex. [^12] For general references on uniform posterior-separable costs, see Frankel and Kamenica (2019), Denti et al. (2022), Denti (2022), and Caplin et al. (2022).
The buyer’s optimization problem is
$u(x,\theta)=\max_{0\leq x^{\prime}\leq x}\left[2\theta x^{\prime}-C(x^{\prime})\right].$
Defining $v(x^{\prime})\equiv-C(x^{\prime})$ and rescaling the type parameter to absorb the coefficient yields
$u(x,\theta)=\max_{0\leq x^{\prime}\leq x}\left[v(x^{\prime})+\theta x^{\prime}\right],$
where $v(\cdot)$ is strictly concave. In this setting, free disposability corresponds to the buyer’s option to conduct a coarser experiment than the dataset permits, and zero marginal cost reflects that, once data is aggregated, providing higher-precision datasets entails no additional production cost.
elation to Literature
The framework of Coasian bargaining was introduced by Coase (1972) and subsequently developed by Fudenberg et al. (1985), Gul et al. (1986), and Ausubel and Deneckere (1989b). This literature shows that when both parties are patient, the lack of commitment eliminates the seller’s bargaining power. The approximate equilibrium outcome is that the buyer pays a price equal to their lowest possible valuation, and trade occurs almost immediately. This result is known as the Coase conjecture.
The literature provides several interpretations of the forces underlying Coasian dynamics. One interpretation is that limited commitment, combined with one-sided private information, drives the market toward efficiency. In the patient limit, if the market clears almost immediately, no buyer type experiences delay. Consequently, the discounted allocation probability for every type approaches one. This outcome is allocatively efficient because, when all types have strictly positive valuations, maximizing social surplus requires trading with every type with probability one.
Nava and Schiraldi (2019) offer an alternative perspective based on market clearing. They show that in settings with multiple varieties (exogenous quality differentiation), the seller’s equilibrium profit is bounded below by the optimal market-clearing profit, converging to this bound in the patient limit across all weak-Markov equilibria. By providing a menu of varieties within a single period, the seller retains the ability to distort allocations through intratemporal price discrimination. This capacity persists even in the patient limit; consequently, full market efficiency is generally not attained.
Nava and Schiraldi (2019) subsumes prior findings. It aligns with the single-good models of Fudenberg et al. (1985) and Gul et al. (1986), where the optimal market-clearing profit coincides with the efficient outcome in terms of allocation probability. At the same time, it accommodates instances of “Coasian failure,” such as the multi-variety bargaining models of Wang (1998), Hahn (2006), and Mensch (2017), as well as the outside-option framework of Board and Pycia (2014). In the former group, the static optimal menu clears the market, so intratemporal price discrimination restores monopoly power despite the absence of intertemporal commitment. In Board and Pycia (2014), a positive outside option facilitates market clearing at the end of the initial period, implying that posting the optimal static price allows the seller to achieve the commitment payoff, defined as the valuation net of the outside option.
This highlights a restriction implicit in the models of Fudenberg et al. (1985) and Gul et al. (1986): the seller offers a single price in each period. Without quality differentiation, this constraint is less restrictive because the Myersonian optimal mechanism is a posted price. In settings with quality differentiation (e.g., Mussa and Rosen (1978)), however, the restriction becomes substantive. It precludes intratemporal price discrimination and limits the seller to intertemporal price discrimination.
Our results indicate that within the Coasian bargaining framework, optimal market-clearing profit and market efficiency (or the Folk Theorem) constitute distinct economic forces. The first, identified by Nava and Schiraldi (2019), derives from intratemporal price discrimination. The second, which is the focus of this analysis, stems from intertemporal price discrimination.
Standard Model
Fudenberg et al. (1985), Gul et al. (1986), Ausubel and Deneckere (1989b)
(No Quality Differentiation)
This Paper
(Quality Differentiation)
Result: Folk Theorem
Nava and Schiraldi (2019)
(Quality Differentiation)
Result: Optimal Market-Clearing Profit
Focus: intertemporalPrice DiscriminationFocus: intratemporalPrice Discrimination
Figure 1: Conceptual Relationship to the Standard Coasian Model
Several other papers in the Coasian bargaining literature relate to our work. First, Doval and Skreta (2024) show that in the single-quality setting, the Coase conjecture holds even when the seller can offer arbitrary mechanisms within each period. The distinction lies in the modeling of differentiation: Doval and Skreta (2024) rely on endogenous allocation probabilities, whereas Nava and Schiraldi (2019) employ exogenous quality differentiation. This difference implies that an allocation probability strictly less than one in Doval and Skreta (2024) does not necessarily remove the buyer type from the market, whereas under quality differentiation, any type accepting an offer leaves the market with probability one because allocation operates along the quality dimension rather than probability. Consequently, although Doval and Skreta (2024) allow for intratemporal discrimination, their framework does not yield outcomes analogous to Wang (1998), Hahn (2006), or Mensch (2017), because the optimal market-clearing profit in their setting remains the lowest possible buyer valuation. Thus, the finding in Doval and Skreta (2024) is consistent with the general principle established by Nava and Schiraldi (2019).
Second, Ali et al. (2023) analyze sequential bargaining with a veto player whose single-peaked preferences allow the proposer to “leapfrog.” In their setting, this preference structure (single crossing in both directions) enables the proposer to clear the market from the bottom up, thereby approximating the commitment payoff. By comparison, in our model a higher-type buyer always derives strictly greater utility from a given offer. This monotonicity ensures that the lowest type remains active until the game concludes, precluding the bottom-up clearing dynamics observed in their environment.
Finally, a distinct strand of literature examines the Coase conjecture within contract-renegotiation frameworks (e.g., Strulovici, 2017; Maestri, 2017). These studies show that as frictions vanish, equilibrium outcomes converge to the efficient allocation. Our framework differs from this literature in its strategy space: we model a sequence of take-it-or-leave-it offers, whereas these papers study the renegotiation of existing contracts.
Beyond Coasian bargaining, our work contributes to the literature on nonlinear pricing (e.g., Mussa and Rosen, 1978; Maskin and Riley, 1984; Wilson, 1993) and mechanism design with ex-post moral hazard (e.g., Laffont and Tirole, 1986; Carbajal and Ely, 2013; Strausz, 2017; Gershkov et al., 2021). Regarding recent contributions, Yang (2022) characterize revenue-maximizing mechanisms for data brokers selling information to producers for price discrimination. Corrao et al. (2023) analyze nonlinear pricing for goods whose usage generates revenue but is subject to free disposal. To our knowledge, this paper is the first to explicitly analyze dynamic screening and allocation under ex-post moral hazard in a limited-commitment setting.
Lastly, our work relates to the literature on reputational bargaining, initiated by Abreu and Gul (2000). While that literature rationalizes delay and inefficiency through the presence of behavioral types, we do not model commitment as an exogenous characteristic. Instead, we identify reputational effects that arise endogenously from the structural properties of digital goods.
onclusion
Our results show how specific technological features alter the strategic structure of dynamic monopoly. When allocative efficiency is set-valued rather than single-valued such as in digital goods markets, Coasian forces do not generally select a unique outcome. Given the indeterminacy, realized outcomes may depend on norms, reputation, or seller-specific characteristics rather than solely on fundamentals such as demand, buyer valuations, or production costs.
This implication is consistent with observed patterns in the B2B digital goods sector, as illustrated by Broadcom’s acquisition of VMware. VMware’s core product, vSphere, is a server virtualization platform, a software layer that partitions a physical server into multiple virtual machines. As documented, Broadcom terminated existing licenses and required customers to renegotiate, resulting in revised terms, with renewal prices increasing by 800% and, in some cases, exceeding 1,500%  [^13] See https://www.ciodive.com/news/broadcom-vmware-lock-in-cost-disruption/718281/.. Although the seller changed from VMware to Broadcom, the product itself did not change.
This raises the relevance of regulatory oversight. Although optimal policy design lies beyond the scope of this paper, the analysis suggests that market forces alone may not generate consistent or efficient outcomes.
Appendix A Proofs
A.1 Proofs of the Results in Section 3
Proof of Lemma 3.
Following standard arguments in static mechanism design, we first solve a relaxed version of the seller’s problem. We then verify that the resulting solution satisfies the original constraints, which establishes optimality.
  1. Step 1.
    Relaxed Problem. Define the information rent function for each type $\theta$ as
    $U(\theta)=\max_{\theta^{\prime}}\left\{q(\theta^{\prime})\,\mathbb{E}\bigl[u(x(\theta^{\prime}),\theta)\bigr]-p(\theta^{\prime})\right\}.$
    By the Dominated Convergence Theorem, the derivative of expected utility with respect to $\theta$ satisfies
    $\frac{\partial\mathbb{E}[u(x(\theta),\theta)]}{\partial\theta}=\mathbb{E}\!\left[\frac{\partial u(x(\theta),\theta)}{\partial\theta}\right].$
    Applying the Envelope Theorem yields
    $U^{\prime}(\theta)=q(\theta)\frac{\partial\mathbb{E}[u(x(\theta),\theta)]}{\partial\theta}=q(\theta)\mathbb{E}\bigl[\min\{x^{e}(\theta),x(\theta)\}\bigr].$
    It is optimal to set $U(\underline{\theta})=0$, so that the individual rationality constraint binds at the lowest type. The seller’s relaxed problem then reduces to maximizing expected revenue. Applying integration by parts to the double integral term yields
    $\int_{\underline{\theta}}^{\overline{\theta}}p(\theta)\,\mathrm{d}F(\theta)$ $=\int_{\underline{\theta}}^{\overline{\theta}}\Bigg[q(\theta)\mathbb{E}\bigl[v(\min\{x^{e}(\theta),x(\theta)\})\bigr]+\theta\,\mathbb{E}\bigl[\min\{x^{e}(\theta),x(\theta)\}\bigr]$
    $\quad-\int_{\underline{\theta}}^{\theta}q(z)\mathbb{E}\bigl[\min\{x^{e}(z),x(z)\}\bigr]\,\mathrm{d}z\Bigg]f(\theta)\,\mathrm{d}\theta$
    $=\int_{\underline{\theta}}^{\overline{\theta}}q(\theta)\left[\mathbb{E}\bigl[v(\min\{x^{e}(\theta),x(\theta)\})\bigr]+\mathbb{E}\bigl[\min\{x^{e}(\theta),x(\theta)\}\bigr]\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right)\right]f(\theta)\,\mathrm{d}\theta.$
    We define the virtual surplus for type $\theta$ given consumption $x$ as
    $\phi(x,\theta)=v(x)+x\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right).$
    Let $x^{m}(\theta)$ denote the allocation that maximizes the virtual surplus,
    $x^{m}(\theta)=\arg\max_{x\in[0,\overline{x}]}\phi(x,\theta).$
    Since $\phi(\cdot,\theta)$ is strictly concave, $x^{m}(\theta)$ is unique. Subject to the feasibility constraint $x\in[\underline{x},\overline{x}]$, the seller’s optimal choice is the projection of $x^{m}(\theta)$ onto this interval. Thus, the optimal feasible allocation is
[…]
    Given the allocation $x^{*}(\theta)$, the consumption level chosen by type $\theta$ is
    $\min\{x^{e}(\theta),x^{*}(\theta)\}.$
    Assumption 3 ensures that the virtual surplus at this consumption level is nonnegative. In particular,
    $\phi(\min\{x^{e}(\theta),x^{*}(\theta)\},\theta)\geq\phi(x^{e}(\theta),\theta)\geq 0.$
    The first inequality follows from the strict concavity of $\phi(\cdot,\theta)$ and the fact that $\min\{x^{e}(\theta),x^{*}(\theta)\}$ lies between $x^{m}(\theta)$ and $x^{e}(\theta)$, or coincides with one of them. Pointwise maximization therefore implies that it is optimal to set $q(\theta)=1$ for all $\theta$.
  2. Step 2.
    Verification of Monotonicity. By Assumption 2, the virtual valuation
    $J(\theta)=\theta-\frac{1-F(\theta)}{f(\theta)}$
    is strictly increasing. Consequently, the unconstrained maximizer $x^{m}(\theta)$ is non-decreasing in $\theta$. Since projection onto the interval $\underline{x},\overline{x}]$ preserves monotonicity, the candidate allocation $x^{*}(\theta)$ is also non-decreasing. Therefore, the solution to the relaxed problem satisfies the monotonicity condition required for incentive compatibility and is optimal for the original mechanism design problem.
Proof of Corollary 1.
Fix a seller strategy $\sigma$ and a buyer best response $\alpha$, and let $(\sigma,\alpha)$ denote the induced strategy profile. For each type $\theta$, this profile induces a probability measure $\lambda_{\theta}$ on $\underline{x},\overline{x}]\times(\mathbb{N}\cup\{\infty\})$ over allocation–timing outcomes, and a probability measure $\nu_{\theta}$ on $\mathbb{R}\times(\mathbb{N}\cup\{\infty\})$ over payment–timing outcomes. An outcome $(x,p,t)\in[\underline{x},\overline{x}]\times\mathbb{R}\times\mathbb{N}$ indicates that the offer $(x,p)$ is accepted in period $t$, while $t=\infty$ denotes no sale.
For each $t\in\mathbb{N}$, let $\lambda_{\theta}(t)$ denote the measure on $\underline{x},\overline{x}]$ defined by
$\lambda_{\theta}(t)(B):=\lambda_{\theta}(B\times\{t\}),\quad\text{for every Borel set }B\subseteq[\underline{x},\overline{x}].$
Similarly, define $\nu_{\theta}(t)$ as the measure on $\mathbb{R}$ given by
$\nu_{\theta}(t)(B):=\nu_{\theta}(B\times\{t\}),\quad\text{for every Borel set }B\subseteq\mathbb{R}.$
We define the components of the equivalent static mechanism $(q,\tilde{x},p)$ as follows:
$q(\theta)$ $:=\sum_{t=0}^{\infty}\delta^{t}\lambda_{\theta}(t)([\underline{x},\overline{x}]),$
$\tilde{x}(\theta)$ $\sim\frac{1}{q(\theta)}\sum_{t=0}^{\infty}\delta^{t}\lambda_{\theta}(t)\quad\text{(well-defined if }q(\theta)>0\text{)},$
$p(\theta)$ $:=\sum_{t=0}^{\infty}\delta^{t}\int_{\mathbb{R}}p\,\mathrm{d}\nu_{\theta}(t).$
Consequently, for any pair $(\theta,\theta^{\prime})$, the expected payoff in the static setting satisfies
$q(\theta^{\prime})\,\mathbb{E}_{\tilde{x}(\theta^{\prime})}[u(x,\theta)]-p(\theta^{\prime})=\sum_{t=0}^{\infty}\delta^{t}\left(\int_{[\underline{x},\overline{x}]}u(x,\theta)\,\mathrm{d}\lambda_{\theta^{\prime}}(t)-\int_{\mathbb{R}}p\,\mathrm{d}\nu_{\theta^{\prime}}(t)\right).$
The right-hand side equals the expected payoff for type $\theta$ when mimicking type $\theta^{\prime}$ in the dynamic game. Since $\alpha$ is a best response for the buyer, the induced static mechanism satisfies incentive compatibility and individual rationality. By construction, the seller’s expected payoff in the static mechanism coincides with her payoff under $(\sigma,\alpha)$ in the dynamic game. Therefore, every dynamic payoff can be replicated in the static setting, which implies that the optimal static payoff is an upper bound on the seller’s achievable payoffs.
Proof of Lemma 4.
We follow the same logic as in the proof of Lemma 3. Recall that the virtual surplus for type $\theta$ given consumption $x$ is
$\phi(x,\theta)=v(x)+x\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right).$
Let
$x^{m}(\theta)\in\arg\max_{x\in[0,\overline{x}]}\phi(x,\theta)$
denote the unconstrained allocation that maximizes virtual surplus. Since $\phi(\cdot,\theta)$ is strictly concave, $x^{m}(\theta)$ is unique.
We impose the additional constraints that $q(\underline{\theta})=1$ and $x(\underline{\theta})\geq x^{e}(\underline{\theta})$ almost surely. Incentive compatibility then requires that, for every type $\theta$, the condition $\mathbb{E}[x(\theta)]\geq x^{e}(\underline{\theta})$ holds. Combining these restrictions, we define the lower bound on feasible allocations as
$L\equiv\max\{x^{e}(\underline{\theta}),\underline{x}\}.$
Thus, the revenue-maximizing allocation is obtained by projecting the unconstrained maximizer $x^{m}(\theta)$ onto the restricted interval $L,\overline{x}]$. Since the virtual surplus function is strictly concave, randomization is not optimal. The resulting allocation is
$x^{*}(\theta)=\begin{cases}x^{m}(\theta),&\text{if }x^{m}(\theta)\geq L,\\[6.0pt] L,&\text{if }x^{m}(\theta)<L.\end{cases}$
Assumption 3 ensures that the virtual surplus at the chosen consumption level is nonnegative. In particular,
$\phi(\min\{x^{e}(\theta),x^{*}(\theta)\},\theta)\geq\phi(x^{e}(\theta),\theta)\geq 0.$
The first inequality follows from the strict concavity of $\phi(\cdot,\theta)$ and the fact that $x^{*}(\theta)$ lies weakly closer to the unconstrained maximizer $x^{m}(\theta)$ than $x^{e}(\theta)$. Therefore, it is optimal to set $q(\theta)=1$ for all $\theta$.
Finally, by Assumption 2, the virtual valuation is strictly increasing, which implies that $x^{m}(\theta)$ is non-decreasing. Since projection onto $L,\overline{x}]$ preserves monotonicity, $x^{*}(\theta)$ satisfies the incentive compatibility constraint. The solution to the relaxed problem is therefore feasible for the original mechanism design problem and is optimal.
A.2 Proofs of the Results in Section 4
The following two lemmas are standard auxiliary results that establish why immediate market clearing is always feasible for the seller.
Lemma 5.
Suppose $(\sigma,\alpha)$ are equilibrium strategies. At any history $h_{t}$, let $A(h_{t})$ denote the set of active buyer types, that is, those who have not yet purchased, and let $\underline{\theta}(h_{t})=\inf A(h_{t})$. Let $(x_{t},p_{t})\in\sigma(h_{t})$. Then
$u(x_{t},\underline{\theta}(h_{t}))-p_{t}\leq 0.$
Proof.
Assume the contrary. Suppose there exists a history $h_{t}$ and an offer $(x_{t},p_{t})\in\sigma(h_{t})$ such that
$u(x_{t},\underline{\theta}(h_{t}))-p_{t}>0.$
Define
$S=\left\{h:\sup_{(x,p)\in\sigma(h)}\bigl(u(x,\underline{\theta}(h))-p\bigr)>0\right\},$
and
$d_{\sup}=\sup_{h\in H}\sup_{(x,p)\in\sigma(h)}\bigl\{u(x,\underline{\theta}(h))-p\bigr\}.$
Since $S\neq\emptyset$, we have $d_{\sup}>0$. Hence, there exist $h_{t}^{*}\in S$ and $(x_{t}^{*},p_{t}^{*})\in\sigma(h_{t}^{*})$ such that
$u(x_{t}^{*},\underline{\theta}(h_{t}^{*}))-p_{t}^{*}>d_{\sup}-\epsilon$
for some $\epsilon>0$.
We claim that all active buyer types accept any offer $(x_{t}^{*},p_{t})$ with $p_{t}\leq p_{t}^{*}+\epsilon$ at $h_{t}^{*}$. Indeed, such an offer yields each type at least
$u(x_{t}^{*},\underline{\theta}(h_{t}^{*}))-(p_{t}^{*}+\epsilon)>d_{\sup}-2\epsilon,$
since higher types can mimic $\underline{\theta}(h_{t}^{*})$ and obtain weakly higher utility. If instead a nonempty set of types waits, then for any continuation history $h_{s}^{*}\supset h_{t}^{*}$,
$\sup_{(x_{s}^{*},p_{s}^{*})\in\sigma(h_{s}^{*})}\bigl(u(x_{s}^{*},\underline{\theta}(h_{s}^{*}))-p_{s}^{*}\bigr)\leq d_{\sup}.$
By continuity of $u(\cdot,\cdot)$ in $\theta$, there exists $\theta_{s}^{*}\in A(h_{s}^{*})$ sufficiently close to $\underline{\theta}(h_{s}^{*})$ such that
$\sup_{(x_{s}^{*},p_{s}^{*})\in\sigma(h_{s}^{*})}\bigl(u(x_{s}^{*},\theta_{s}^{*})-p_{s}^{*}\bigr)\leq d_{\sup}+\epsilon.$
If $\epsilon$ is chosen such that
$d_{\sup}-2\epsilon>\delta(d_{\sup}+\epsilon),$
then type $\theta_{s}^{*}$ strictly prefers to purchase immediately at $h_{t}^{*}$, contradicting the optimality of waiting.
If all active types accept at $p_{t}\leq p_{t}^{*}+\epsilon$, then the seller can profitably increase the price slightly while still clearing the market, which contradicts the optimality of $\sigma$. Hence, along the equilibrium path, it must hold that
$u(x_{t},\underline{\theta}(h_{t}))-p_{t}\leq 0$
for all $(x_{t},p_{t})\in\sigma(h_{t})$.
Lemma 6.
Any offer $(x_{t},p_{t})$ satisfying (i) $u(x_{t},\underline{\theta})-p_{t}>0$, or (ii) $u(x_{t},\underline{\theta})-p_{t}=0$ with $x_{t}\geq x^{e}(\underline{\theta})$, clears the market immediately, assuming that type $\underline{\theta}$ breaks indifference in favor of the seller.
Proof.
  1. 1.
    If $u(x_{t},\underline{\theta})-p_{t}>0$, then type $\underline{\theta}$ purchases immediately. By Lemma 5, she anticipates no strictly positive surplus in any future period.
    Suppose instead that at history $h_{t+1}$ there remain active buyers, that is, $A(h_{t+1})\neq\emptyset$. Consider a sequence $\{\theta_{i}\}\subset A(h_{t+1})$ with $\theta_{i}\downarrow\inf A(h_{t+1})$. By Lemma 5, waiting yields nonpositive surplus for type $\inf A(h_{t+1})$. By continuity, there therefore exists some type $\theta\in A(h_{t+1})$ whose expected surplus from waiting is arbitrarily close to zero. Such a buyer should instead have purchased at $h_{t}$, since doing so would have yielded at least
    $u(x_{t},\inf A(h_{t+1}))-p_{t}\geq u(x_{t},\underline{\theta})-p_{t}>0.$
  2. 2.
    Now suppose that $u(x_{t},\underline{\theta})-p_{t}=0$ and $x_{t}\geq x^{e}(\underline{\theta})$. Again, suppose to the contrary that $A(h_{t+1})\neq\emptyset$. If $\inf A(h_{t+1})>\underline{\theta}$, the argument above applies directly. If $\inf A(h_{t+1})=\underline{\theta}$, consider a sequence $\{\theta_{i}\}\subset A(h_{t+1})$ with $\theta_{i}\downarrow\underline{\theta}$. By waiting, type $\theta_{i}$ obtains at most
    $\delta\bigl[u(x^{e}(\theta_{i}),\theta_{i})-u(x^{e}(\underline{\theta}),\underline{\theta})\bigr]=\delta\Big(u(x^{e}(\theta_{i}),\theta_{i})-u(x^{e}(\underline{\theta}),\theta_{i})+u(x^{e}(\underline{\theta}),\theta_{i})-u(x^{e}(\underline{\theta}),\underline{\theta})\Big).$
    Observe that
    $u(x^{e}(\theta_{i}),\theta_{i})-u(x^{e}(\underline{\theta}),\theta_{i})$ $=\int_{x^{e}(\underline{\theta})}^{x^{e}(\theta_{i})}\bigl(v^{\prime}(z)+\theta_{i}\bigr)\,\mathrm{d}z$
    $\leq\bigl[x^{e}(\theta_{i})-x^{e}(\underline{\theta})\bigr]\cdot\bigl[v^{\prime}(x^{e}(\underline{\theta}))+\theta_{i}\bigr]$
    $=\bigl[x^{e}(\theta_{i})-x^{e}(\underline{\theta})\bigr]\cdot(\theta_{i}-\underline{\theta}),$
    where the inequality follows from the fact that $v^{\prime}(\cdot)$ is decreasing, and the final equality uses the first-order condition $v^{\prime}(x^{e}(\underline{\theta}))+\underline{\theta}=0$. Therefore,
    $\delta\bigl[u(x^{e}(\theta_{i}),\theta_{i})-u(x^{e}(\underline{\theta}),\underline{\theta})\bigr]$ $\leq\delta\Bigl[\bigl(x^{e}(\theta_{i})-x^{e}(\underline{\theta})\bigr)(\theta_{i}-\underline{\theta})+u(x^{e}(\underline{\theta}),\theta_{i})-u(x^{e}(\underline{\theta}),\underline{\theta})\Bigr]$
    $\leq\delta\Bigl[\bigl(x^{e}(\theta_{i})-x^{e}(\underline{\theta})\bigr)(\theta_{i}-\underline{\theta})+(\theta_{i}-\underline{\theta})\,x^{e}(\underline{\theta})\Bigr].$
    As $\theta_{i}\to\underline{\theta}$, we have $x^{e}(\theta_{i})-x^{e}(\underline{\theta})\to 0$. For $\delta<1$ and $x^{e}(\underline{\theta})>0$, it follows that for $\theta_{i}$ sufficiently close to $\underline{\theta}$,
    $\delta\bigl[u(x^{e}(\theta_{i}),\theta_{i})-u(x^{e}(\underline{\theta}),\underline{\theta})\bigr]<(\theta_{i}-\underline{\theta})\,x^{e}(\underline{\theta})=u(x^{e}(\underline{\theta}),\theta_{i})-u(x^{e}(\underline{\theta}),\underline{\theta}).$
    The right-hand side is the payoff from purchasing immediately by mimicking type $\underline{\theta}$ when the offered allocation satisfies $x_{t}\geq x^{e}(\underline{\theta})$. Hence, type $\theta_{i}$ strictly prefers to purchase immediately, which contradicts sequential rationality.
Proof of Theorem 1.
We begin with a preliminary lemma that characterizes the optimal uniform price when all types consume their efficient allocations.
Lemma 7.
Let $F_{\theta^{\prime}}(\cdot)$ denote the buyer type distribution conditional on the interval $\underline{\theta},\theta^{\prime}]$ for some $\theta^{\prime}\in\Theta$. If each type $\theta$ consumes the efficient quantity $x^{e}(\theta)$ and the seller is restricted to charging a single uniform price $p$, then the revenue-maximizing price is
$p=u_{\min}.$
Proof.
Given the efficient allocation $x^{e}(\theta)$, the buyer’s gross utility is
$v(x^{e}(\theta))+\theta\,x^{e}(\theta).$
By Assumption 3, the virtual surplus satisfies
$v(x^{e}(\theta))+\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right)x^{e}(\theta)\;\geq\;0$
for all $\theta$. Consequently, the seller optimally sets the allocation probability $q(\theta)=1$ for all $\theta$ and charges the market-clearing price $u_{\min}$.
Next, consider the conditional distributions $F_{\theta^{\prime}}(\cdot)$ and $f_{\theta^{\prime}}(\cdot)$. For all $\theta\leq\theta^{\prime}$, we have
$\theta-\frac{1-F(\theta)}{f(\theta)}\;\leq\;\theta-\frac{1-F_{\theta^{\prime}}(\theta)}{f_{\theta^{\prime}}(\theta)}.$
Thus, the virtual surplus increases pointwise under the conditional distribution. It follows that $q(\theta)=1$ and $p=u_{\min}$ remain optimal.
We now turn to the proof of the theorem. The first step is to define the cutoff type $\theta(\delta,x,p)$ for a given offer $(x,p)$.
  1. 1.
    If $u(x,\underline{\theta})-p\geq 0$, we set $\theta(\delta,x,p)=\underline{\theta}$. In this case, every active type purchases immediately.
  2. 2.
    If $u(x,\underline{\theta})-p<0$, let $\theta(\delta,x,p)$ denote the unique solution, if it exists, to the indifference condition
    $u(x,\theta)-p=\delta\big(u(\max\{x,x^{e}(\underline{\theta})\},\theta)-u_{\min}\big).$
    By construction, type $\theta(\delta,x,p)$ is indifferent between purchasing immediately at price $p$ and waiting to purchase the allocation $\max\{x,x^{e}(\underline{\theta})\}$ in the subsequent period at price $u_{\min}$. If a solution exists, it is unique by the single-crossing property of the utility function. Although the net utility $u(x,\theta)-p$ need not satisfy single crossing in $\theta$ for a fixed offer $(x,p)$ in general, our construction of the continuation allocation $\max\{x,x^{e}(\underline{\theta})\}$ ensures that the indifference condition exhibits single crossing. In particular, the difference between the current-period payoff and the discounted continuation payoff crosses zero at most once in $\theta$, which guarantees uniqueness of the cutoff type.
    Consider two cases. First, if $x\geq x^{e}(\underline{\theta})$, the result follows immediately because the allocation levels coincide across periods and the continuation payoff is a discounted version of the current payoff. Second, if $x<x^{e}(\underline{\theta})$, the continuation payoff $\delta\big(u(x^{e}(\underline{\theta}),\theta)-u_{\min}\big)$ is linear in $\theta$. This linearity holds because all types $\theta\geq\underline{\theta}$ optimally consume at least $x^{e}(\underline{\theta})$ and are therefore constrained to consume exactly the offered amount. Since the current-period utility $u(x,\theta)-p$ is convex in $\theta$ and strictly negative at $\underline{\theta}$, it can intersect the linear continuation payoff from below at most once.
  3. 3.
    If $u(x,\underline{\theta})-p<0$ and the indifference equation admits no solution, we set $\theta(\delta,x,p)=\overline{\theta}$. In this case, every type chooses to wait.
We consider the following candidate equilibrium strategies:
  1. 1.
    Seller: In each period $t$, offer $(x_{t},p_{t})$ with $x_{t}=\max\{x_{t-1},x^{e}(\underline{\theta})\}$ and $p_{t}=u_{\min}$, starting from $x_{-1}=\overline{x}$. Denote this strategy by $\sigma$.
  2. 2.
    Buyer: In period $t$, all types $\theta\geq\theta(\delta,x_{t},p_{t})$ accept immediately, while types $\theta<\theta(\delta,x_{t},p_{t})$ wait. Denote this strategy by $\alpha$.
Given any offer $(x_{t},p_{t})$, if $u(x_{t},\underline{\theta})-p_{t}<0$, the buyer anticipates the continuation offer $(\max\{x_{t},x^{e}(\underline{\theta})\},u_{\min})$ in the subsequent period. Given this expectation, $\alpha$ is a best response. Conversely, if $u(x_{t},\underline{\theta})-p_{t}\geq 0$, the market clears immediately.
We now show that $\sigma$ is optimal for the seller. First, note that any deviation satisfying $u(x,\underline{\theta})-p\geq 0$ cannot be strictly optimal, since $u_{\min}$ is the maximum attainable market-clearing profit.
Suppose, for a contradiction, that the seller has a strictly profitable deviation $\sigma^{\prime}$ at some history $h_{t}$. Because $\delta<1$ and payoffs are bounded, the One-Shot Deviation Principle implies that there exists a finite $n$ such that deviating in periods $t,\dots,t+n-1$ and reverting to $\sigma$ at period $t+n$ remains strictly profitable.
Consider the final deviation period t+n1t+n-1, in which the seller offers $(x_{t+n-1},p_{t+n-1})$. Under the reversion strategy, the seller clears the remaining market in period $t+n$ with the offer $(\max\{x_{t+n-1},x^{e}(\underline{\theta})\},u_{\min})$.
Construct an alternative offer $(x^{\prime}_{t+n-1},p^{\prime}_{t+n-1})$ defined by
$x^{\prime}_{t+n-1}=\max\{x^{e}(\theta(\delta,x_{t+n-1},p_{t+n-1})),\underline{x}\},$
and choose $p^{\prime}_{t+n-1}$ such that
$u(x^{\prime}_{t+n-1},\theta_{t+n-1})-p^{\prime}_{t+n-1}=\delta\Big(u(\max\{x^{\prime}_{t+n-1},x^{e}(\underline{\theta})\},\theta_{t+n-1})-u_{\min}\Big).$
This modification preserves the cutoff type $\theta_{t+n-1}$ and yields a weakly higher price. Since the allocation is efficient, we have
$(1-\delta)u(x^{\prime}_{t+n-1},\theta_{t+n-1})\;\geq\;(1-\delta)u(x_{t+n-1},\theta_{t+n-1}).$
It follows that
$p^{\prime}_{t+n-1}\geq p_{t+n-1}.$
Consequently, the seller weakly improves her payoff by offering $(x^{\prime}_{t+n-1},p^{\prime}_{t+n-1})$ in period t+n1t+n-1 and then clearing the market with $(\max\{x^{\prime}_{t+n-1},x^{e}(\underline{\theta})\},u_{\min})$ in period $t+n$.
However, by Lemma 7, immediate market clearing in period t+n1t+n-1 at price $p=u_{\min}$ yields a strictly higher payoff than delaying market clearing to period $t+n$. Thus, the deviation is strictly dominated by immediate market clearing in period t+n1t+n-1. By backward induction, the same argument applies to all preceding deviation periods. We therefore conclude that no profitable deviation exists, and $(\sigma,\alpha)$ constitutes an equilibrium for all $0\leq\delta<1$.
Proof of Theorem 2.
We proceed in two steps.
  1. Step 1.
    Construction of the high-payoff equilibrium. We show that for any $\epsilon>0$, there exists a discount factor $\underline{\delta}$ such that for all $\delta\geq\underline{\delta}$, one can construct an equilibrium with payoff at least $\pi^{e}(F)-\epsilon$.
    Recall from Lemma 4 that the optimal mechanism is given by
    $x^{*}(\theta)=\begin{cases}x^{m}(\theta),&\text{if }x^{m}(\theta)>L,\\[6.0pt] L,&\text{if }x^{m}(\theta)\leq L,\end{cases}$
    where $L\equiv\max\{x^{e}(\underline{\theta}),\underline{x}\}$. Let $\theta^{*}$ denote the unique type such that $x^{m}(\theta^{*})=L$.
    Construct a partition $\{\theta_{0},\dots,\theta_{n}\}$ of the type space $\underline{\theta},\overline{\theta}]$ such that
    $\overline{\theta}=\theta_{0}>\theta_{1}>\dots>\theta_{n}=\theta^{*},$
    with mesh size $\max_{i}|\theta_{i-1}-\theta_{i}|$ approaching zero as $n\to\infty$. Assign allocations $x_{1},\dots,x_{n}$ such that
    $x_{i}=x^{m}(\theta_{i})\ \text{for }i=1,\dots,n-1,\qquad x_{n}=L.$
    Set the final price $p_{n}=u_{\min}$. Define the remaining prices $p_{1},\dots,p_{n-1}$ recursively by backward induction such that for each cutoff $\theta_{i}$,
    $u(x_{i},\theta_{i})-p_{i}=\delta\,[u(x_{i+1},\theta_{i})-p_{i+1}].$
    Consider the following strategy profile:
    1. (a)
      On-path: In period t=i1t=i-1 for $i=1,\dots,n-1$, the seller posts $(x_{i},p_{i})$, and all active buyers with types $\theta\in[\theta_{i},\theta_{i-1}]$ purchase immediately. In the final period t=n1t=n-1, the seller posts $(x_{n},p_{n})=(L,u_{\min})$ and clears the market on $\theta_{n},\theta_{n-1}]\cup[\underline{\theta},\theta_{n}]$.
    2. (b)
      Off-path: Any deviation by the seller triggers an immediate reversion to the Coasian equilibrium strategies described in Theorem 1.
    The buyer’s incentive compatibility on the equilibrium path holds by construction of the prices $p_{i}$. The seller is also incentive compatible, since any deviation yields a payoff of at most $u_{\min}$, which is weakly less than the payoff from the constructed sequence at every history on the equilibrium path.
    To evaluate the payoff, define the step function $x_{n}(\theta):=x_{i}$ for $\theta\in(\theta_{i},\theta_{i-1}]$. The seller’s total discounted payoff is
    $\sum_{i=1}^{n}\delta^{i-1}\big(F(\theta_{i-1})-F(\theta_{i})\big)p_{i}\\ =\int_{\underline{\theta}}^{\overline{\theta}}\delta^{t(\theta)}\Bigg[v\big(\min\{x_{n}(\theta),x^{e}(\theta)\}\big)+\min\{x_{n}(\theta),x^{e}(\theta)\}\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right)\Bigg]f(\theta)\,\mathrm{d}\theta,$
    where $t(\theta)=i-1$ whenever $\theta\in(\theta_{i},\theta_{i-1}]$.
    We now apply a double-limit argument. First, as $n\to\infty$, the step function converges pointwise, that is, $\min\{x_{n}(\theta),x^{e}(\theta)\}\to\min\{x^{*}(\theta),x^{e}(\theta)\}$. By the Dominated Convergence Theorem, there exists an integer $N$ such that for all $n\geq N$, the distance between the static integrals satisfies
    $\left|\pi^{e}(F)-\int_{\underline{\theta}}^{\overline{\theta}}\left[v\big(\min\{x_{n}(\theta),x^{e}(\theta)\}\big)+\min\{x_{n}(\theta),x^{e}(\theta)\}\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right)\right]f(\theta)\,\mathrm{d}\theta\right|\leq\frac{\epsilon}{2}.$
    Second, fix such an $n\geq N$. Since $\delta^{t(\theta)}\to 1$ as $\delta\to 1$ for fixed finite $n$, there exists a threshold $\underline{\delta}(n)$ such that for all $\delta>\underline{\delta}(n)$:
    $\Bigg|\int_{\underline{\theta}}^{\overline{\theta}}\delta^{t(\theta)}\left[v\big(\min\{x_{n}(\theta),x^{e}(\theta)\}\big)+\min\{x_{n}(\theta),x^{e}(\theta)\}\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right)\right]f(\theta)\,\mathrm{d}\theta\\ -\int_{\underline{\theta}}^{\overline{\theta}}\left[v\big(\min\{x_{n}(\theta),x^{e}(\theta)\}\big)+\min\{x_{n}(\theta),x^{e}(\theta)\}\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right)\right]f(\theta)\,\mathrm{d}\theta\Bigg|\leq\frac{\epsilon}{2}.$
    Combining these two inequalities using the triangle inequality yields a total payoff of at least $\pi^{e}(F)-\epsilon$.
  2. Step 2.
    Filling the Continuum. We show that any payoff in the interval $u_{\min},\,\pi^{e}(F)-\epsilon]$ is sustainable. Let $L\equiv\max\{x^{e}(\underline{\theta}),\underline{x}\}$. For a parameter $s\in[0,1]$, define the convex combination of allocations by
    $x_{i}(s)=sx_{i}+(1-s)L,\quad\text{for }i=1,\dots,n.$
    Set the terminal price $p_{n}(s)=u_{\min}$. Define the remaining prices $p_{1}(s),\dots,p_{n-1}(s)$ recursively such that
    $u(x_{i}(s),\theta_{i})-p_{i}(s)=\delta\,[\,u(x_{i+1}(s),\theta_{i})-p_{i+1}(s)\,].$
    Define the step function $x_{n}(\theta;s):=x_{i}(s)$ for $\theta\in(\theta_{i},\theta_{i-1}]$. The resulting total discounted payoff is
    $\Pi(s)$ $=\sum_{i=1}^{n}\delta^{i-1}\big(F(\theta_{i-1})-F(\theta_{i})\big)p_{i}(s)$
    $=\int_{\underline{\theta}}^{\overline{\theta}}\delta^{t(\theta)}\Bigg[v\big(\min\{x_{n}(\theta;s),x^{e}(\theta)\}\big)$
    $\quad+\min\{x_{n}(\theta;s),x^{e}(\theta)\}\left(\theta-\frac{1-F(\theta)}{f(\theta)}\right)\Bigg]f(\theta)\,\mathrm{d}\theta.$
    At $s=1$, the construction coincides with the high-payoff equilibrium from Step 1, so $\Pi(1)\geq\pi^{e}(F)-\epsilon$. At $s=0$, the allocations satisfy $x_{i}(0)=L$ for all $i$. Off-path deviations trigger an immediate reversion to the Coasian equilibrium, which preserves incentive compatibility for the seller.
    A subtlety arises because $\delta$ is fixed. As the parameter $s$ varies, the on-path payoff, which involves delay, may be dominated by the immediate market-clearing payoff $u_{\min}$. We address this issue by adjusting the horizon. Whenever the on-path continuation payoff equals $u_{\min}$, we reduce the number of periods from $n$ to n1n-1. Under this adjustment, as $s$ approaches 0, the horizon decreases to $n=1$. Given the terminal condition $p_{n}=u_{\min}$, this recursive structure implies that the seller’s payoff converges to $\Pi(0)=u_{\min}$.
    By continuity of the integral with respect to the parameter $s$, $\Pi(s)$ is continuous on $[0,1]$. Therefore, by the Intermediate Value Theorem, for any target payoff $\pi\in[u_{\min},\,\Pi(1)]$, there exists $s^{*}\in[0,1]$ such that $\Pi(s^{*})=\pi$.
Proof of Corollary 4.
The equilibrium construction follows directly from the proof of Theorem 2. To satisfy Definition 2, it suffices to specify the buyer’s strategy for offers $(x_{i},p_{i})$ that occur on the equilibrium path for $i=1,\dots,n$.
Consider the following buyer strategy. Upon observing any offer $(x_{i},p_{i})$ on the equilibrium path, all active types $\theta\geq\theta_{i}$ accept immediately. This specification conditions acceptance solely on the current offer $(x_{i},p_{i})$ and therefore satisfies the on-path Markov property. Following such an offer, play proceeds according to the equilibrium dynamics described in Theorem 2, starting from stage $i+1$.
We show that for $\delta$ sufficiently close to $1$, the seller has no incentive to deviate from the prescribed offer $(x_{i},p_{i})$ to a subsequent on-path offer $(x_{j},p_{j})$ with $j>i$. By construction, the indifference condition implies
$u(x_{i},\theta_{i})-p_{i}=\delta\big[u(x_{i+1},\theta_{i})-p_{i+1}\big].$
For a fixed grid size $n$, the allocation differences $|x_{i+1}-x_{i}|$ are bounded below by some $\epsilon_{1}>0$. Consequently, the corresponding price differences $p_{i}-p_{i+1}$ are bounded below by a constant $\epsilon_{2}>0$. Since the density $f(\cdot)$ is bounded below by $m>0$, the immediate revenue loss incurred on the interval $\theta_{i},\theta_{i-1}]$ by accelerating sales is at least
$m\cdot(\theta_{i-1}-\theta_{i})\cdot(p_{i}-p_{i+1}),$
which is strictly positive and bounded away from zero as $\delta\to 1$. Conversely, the gain from accelerating sales is of order $(1-\delta)K$, where $K$ is a bounded constant representing the future surplus to be collected. This term converges to zero as $\delta\to 1$. Therefore, for $\delta$ sufficiently close to $1$, the potential gain from deviation is strictly dominated by the immediate revenue loss. Under this condition, the seller has no incentive to deviate, which completes the argument.
A.3 Proof of the Results in Section 5
Regardless of the presence of free disposability, the following lemma applies to the discrete-type model with discount factor $\delta\in[0,1)$.
Lemma 8 (Finite-Time Market Clearing).
In any equilibrium, at any history $h_{t}\in H_{t}$, the market clears in finite time $T(\delta)$ along the equilibrium path.
Proof of Lemma 8.
Let $P(h_{t})$ denote the measure of the remaining mass of buyers at history $h_{t}$. We show that for any equilibrium $\{\sigma,\alpha\}$ and any history $h_{t}\in H_{t}$ at which the market has not yet cleared, there exist a positive integer $\kappa(\delta)$ and a constant $k\in(0,1)$ such that
$P(h_{t+\kappa(\delta)})\;\leq\;k\,P(h_{t}).$
Let $\overline{u}(\theta)=v(x^{e}(\theta))+\theta x^{e}(\theta)$. At history $h_{t}$, the seller can guarantee a payoff of at least
$P(h_{t})\,u_{\min}$
by clearing the market immediately. Suppose, for contradiction, that after $\kappa(\delta)$ periods, a proportion $k$ of the mass $P(h_{t})$ remains active. The seller’s total expected profit is then bounded above by
$(1-k)P(h_{t})\,\overline{u}(\overline{\theta})\;+\;\delta^{\kappa(\delta)}k\,P(h_{t})\,\overline{u}(\overline{\theta}).$
Sequential rationality therefore requires
$\big[(1-k)+\delta^{\kappa(\delta)}k\big]\,\overline{u}(\overline{\theta})\;\geq\;u_{\min}>0.$
Thus, the remaining mass must contract by at least the factor $k$ within $\kappa(\delta)$ periods.
To complete the proof, note that the prior distribution contains a point mass at $\underline{\theta}$. Since the remaining mass $P(h_{t})$ cannot decay geometrically below the mass of this atom while the atom remains active, the process must eventually reach a state in which the remaining mass consists predominantly of type $\underline{\theta}$. At that point, the seller clears the market immediately.
Proof of Proposition 1.
By Lemma 8, we apply backward induction.
Since $\sum_{i=1}^{n}q_{i}=1$, the state can be represented by a variable $q\in(0,1]$, which corresponds to the remaining mass of buyers. Let $\theta(q)$ denote the left-continuous function mapping the state $q$ to the corresponding buyer type. At the beginning of period $t$, we denote the state by $q^{t}$.
The structure of the proof follows the logic developed in Gul et al. (1986) and Deneckere and Liang (2006).
Consider first the case in which $q^{t}\leq q_{2}$. In this case, $\theta(q^{t})=\theta_{2}$. The seller’s optimal strategy is therefore to clear the market immediately by offering
$x_{t}=x^{e}(\theta_{2}),\qquad p_{t}=w(x_{t},\theta_{2}).$
Now consider an arbitrary period $t$ and a state $q^{t}$ satisfying $q_{2}<q^{t}\leq 1$. In the final period $t=T(\delta)$, the seller clears the market by offering $(x_{t},p_{t})$ such that
$x_{t}=x^{e}(\theta_{2}),\qquad p_{t}=w(x_{t},\theta_{2}).$
Next, consider the penultimate period $t=T(\delta)-1$. Suppose the seller posts an offer $(x_{t},p_{t})$. We show that the resulting buyer behavior is uniquely determined. First, if
$w(x_{t},\theta_{2})\geq p_{t},$
then the market clears immediately. Otherwise, type $\theta_{2}$ rejects, and type $\theta_{1}$ compares the current payoff with the continuation value. If
$w(x_{t},\theta_{1})-p_{t}>\delta[w(x^{e}(\theta_{2}),\theta_{1})-w(x^{e}(\theta_{2}),\theta_{2})],$
then all buyers of type $\theta_{1}$ purchase immediately; if the inequality is reversed, they wait.
A subtlety arises when this condition holds with equality. Although type $\theta_{1}$ is indifferent, equilibrium requires purchase with probability one. If a strictly positive mass of type $\theta_{1}$ were to wait, the seller could profitably deviate by offering a marginally lower price $p_{t}-\epsilon$, which strictly induces immediate purchase. Therefore, in equilibrium, indifference is resolved in favor of trade.
The seller’s optimization problem is therefore well defined. Given the buyer’s best response, the decision reduces to a comparison between screening type $\theta_{1}$ and clearing the market immediately.
For any $\delta\in[0,1)$, there exists a threshold probability $\bar{q}(\delta)$. If the remaining mass of buyers satisfies $q^{t}>\bar{q}(\delta)$, it is strictly profitable to screen type $\theta_{1}$ by offering $x_{t}=x^{e}(\theta_{1})$. If $q^{t}<\bar{q}(\delta)$, the cost of delaying sales to the larger mass of type $\theta_{2}$ dominates the screening benefit, and the seller strictly prefers to clear the market immediately. When $q^{t}=\bar{q}(\delta)$, the seller is indifferent.
This indifference condition sustains equilibrium existence because type $\theta_{1}$ can randomize so that the posterior mass reaches exactly $\bar{q}(\delta)$. This convexifies the set of buyer continuation payoffs and ensures that an equilibrium exists even following a seller deviation.
This argument explains why the Folk Theorem (Theorem 2) fails in the absence of free disposability. With free disposability, the market-clearing allocation can take any value $x\in[x^{e}(\theta_{2}),\overline{x}]$. This range generates multiplicity in buyer continuation payoffs, which propagates backward to support multiple screening prices for $\theta_{1}$. Without free disposability, the terminal allocation is fixed at $x^{e}(\theta_{2})$, eliminating this source of multiplicity.
Proceeding by backward induction on the horizon, we characterize the equilibrium path. It consists of a sequence of periods in which the seller offers $x^{e}(\theta_{1})$ to screen type $\theta_{1}$ via delay, followed by a final period that clears type $\theta_{2}$. This structure mirrors the standard arguments in Gul et al. (1986) and Deneckere and Liang (2006). By monotone comparative statics (Milgrom and Shannon, 1994), the equilibrium path is unique and deterministic, with the possible exception of randomization in the initial period. If the seller deviates off path, equilibrium existence is guaranteed, though the seller may need to randomize to return to the equilibrium path. Moreover, the essentially unique equilibrium is weak-Markov. As these results are standard in the literature, we omit the detailed derivation.
Lastly, we establish the existence of a uniform bound $T$, independent of $\delta$, on the length of the equilibrium path. This follows from the presence of a point mass at $\theta_{2}$. Accelerating trade, which is feasible under weak-Markov strategies, yields a benefit bounded below by $(1-\delta)L$ for some constant $L>0$. Conversely, the price increment used for screening is constrained by
$p_{t}-p_{t+1}\leq p_{t}-\delta p_{t+1}=(1-\delta)w(x^{e}(\theta_{1}),\theta_{1}).$
Normalizing by the factor $(1-\delta)$, the optimality condition implies that a positive mass of buyers must clear in each period. This requirement yields a bound $T$ that is independent of $\delta$. Consequently, as $\delta\to 1$, the discount factor over the entire horizon, $\delta^{T}$, converges to $1$. The resulting outcome combines efficient allocations with vanishing delay.
Appendix B Additional Results
B.1 Payoff Upper Bound
In this section, we show that $\pi^{e}(F)$ is an upper bound on the seller’s payoff within a broad class of equilibria. As shown above, when $x^{e}(\underline{\theta})\leq\underline{x}$, we have $\pi^{e}(F)=\pi(F)$. This case is comparable to Ausubel and Deneckere (1989b), where the lowest type has zero valuation. In both settings, the lowest type is indifferent across all feasible allocation outcomes, which in the setting of Ausubel and Deneckere (1989b) corresponds to the discounted probability of trade.
We therefore focus on the case in which $x^{e}(\underline{\theta})>\underline{x}$. We begin by introducing the following definitions.
Definition 3 (Horizon and Delay).
Given an equilibrium $\mathcal{E}(\delta)$, let $T(\mathcal{E}(\delta))$ denote the smallest time $t\in\mathbb{N}\cup\{\infty\}$ such that the market clears on the equilibrium path by period $t$. If market clearing never occurs, set $T(\mathcal{E}(\delta))=\infty$.
Definition 4 (On-path Allocation Monotonicity).
An equilibrium satisfies on-path allocation monotonicity if, for every on-path history $h_{t}$, and for any allocations $x_{t}\in\operatorname{supp}\sigma(h_{t})$ and $x_{t-1}\in\operatorname{supp}\sigma(h_{t-1})$ offered on the equilibrium path, we have $x_{t}\leq x_{t-1}$.
We view Definition 4 as a mild restriction in a dynamic adverse-selection environment. The definition constrains behavior only along the equilibrium path. Off-path deviations may be deterred by arbitrary punishment strategies.
Given Definition 4, on-path allocation monotonicity implies the skimming property on the equilibrium path, which is stated in the following corollary.
Corollary 5.
Every equilibrium satisfying on-path allocation monotonicity also satisfies the on-path skimming property. Specifically, on the equilibrium path, if a type $\theta$ weakly prefers purchasing to waiting, then every type $\theta^{\prime}>\theta$ strictly prefers purchasing.
The proof follows directly from the single-crossing property and is omitted. Corollary 5 implies the following lemma.
Lemma 9.
For any given $\delta$, there exists a threshold $\theta^{*}$ such that, in any equilibrium with on-path allocation monotonicity, if the support of the remaining buyer distribution is the interval $\underline{\theta},\theta^{\prime}]$ with $\theta^{\prime}\leq\theta^{*}$, then the seller’s optimal strategy is to clear the market immediately.
Proof of Lemma 9.
Suppose the seller offers a pair $(x_{t},p_{t})$ inducing a cutoff type $\theta$. The seller’s profit is bounded above by
$u(\overline{x},\theta)\,[F(\theta^{\prime})-F(\theta)]+\delta\int_{\underline{\theta}}^{\theta}u(\overline{x},z)f(z)\,\mathrm{d}z,$
since profits are maximized when all purchasing types behave myopically. Differentiating this expression with respect to $\theta$ yields
$\frac{\partial u}{\partial\theta}(\overline{x},\theta)[F(\theta^{\prime})-F(\theta)]-(1-\delta)\,u(\overline{x},\theta)f(\theta).$
By the Envelope Theorem, $\frac{\partial u}{\partial\theta}=\min\{\overline{x},x^{e}(\theta)\}$, so the derivative is bounded above by
$\min\{\overline{x},x^{e}(\theta)\}[F(\theta^{\prime})-F(\theta)]-(1-\delta)\,u(\overline{x},\theta)f(\theta).$
As $\theta^{\prime}\to\underline{\theta}$, the term $F(\theta^{\prime})-F(\theta)$ converges to zero. Moreover, the ratio
$\frac{u(\overline{x},\theta)f(\theta)}{\min\{\overline{x},x^{e}(\theta)\}}$
is uniformly bounded away from zero. Hence, for $\theta^{\prime}$ sufficiently close to $\underline{\theta}$, the derivative is strictly negative for all $\theta\in[\underline{\theta},\theta^{\prime}]$. This implies that the profit bound is strictly decreasing in the cutoff $\theta$ on this interval. Therefore, the maximum is attained at $\theta=\underline{\theta}$, which corresponds to immediate market clearing. ∎
Proposition 2.
Consider a sequence of discount factors $\delta_{n}\to 1$. Let $\mathcal{E}(\delta_{n})$ denote an equilibrium associated with $\delta_{n}$, with seller payoff $\pi(\mathcal{E}(\delta_{n}))$ and on-path market-clearing time $T(\mathcal{E}(\delta_{n}))$. If either
  1. (i)
    each $\mathcal{E}(\delta_{n})$ satisfies on-path allocation monotonicity and Assumption 3 holds, or
  2. (ii)
    delay vanishes asymptotically, that is, $\liminf_{n\to\infty}\delta_{n}^{T(\mathcal{E}(\delta_{n}))}=1$,
then
$\limsup_{n\to\infty}\pi(\mathcal{E}(\delta_{n}))\leq\pi^{e}(F).$
Proposition 2 shows that if the equilibrium exhibits on-path allocation monotonicity, together with Assumption 3, or if delay vanishes asymptotically, then the seller’s payoff in the patient limit is bounded above by $\pi^{e}(F)$.
Proof of Proposition 2.
(i) Assume that each equilibrium in the sequence $(\mathcal{E}(\delta_{n}))$ satisfies on-path allocation monotonicity. The seller’s payoff $\pi(\mathcal{E}(\delta_{n}))$ is bounded above by the value of the following static mechanism design problem:
$\max_{\mathcal{M}}\ \int_{\underline{\theta}}^{\overline{\theta}}p(\theta)f(\theta)\,\mathrm{d}\theta,$
subject to the constraints
$q(\theta)$ $\geq q(\theta^{\prime})\quad\forall\,\theta\geq\theta^{\prime},$
$x(\theta)$ $\geq x(\theta^{\prime})\quad\text{almost everywhere}\quad\forall\,\theta\geq\theta^{\prime},$
$x(\underline{\theta})$ $\geq x^{e}(\underline{\theta})\quad\text{almost everywhere}.$
The first constraint follows from the on-path skimming property implied by on-path allocation monotonicity. The second constraint is on-path allocation monotonicity. The boundary constraint $x(\underline{\theta})\geq x^{e}(\underline{\theta})$ follows from an adaptation of Lemma 8 to the continuum-type setting, using Lemma 9, which shows that the market must clear in finite time and that the market-clearing allocation must be at least $x^{e}(\underline{\theta})$. Under these constraints and Assumption 3, the maximum achievable payoff is $\pi^{e}(F)$. Therefore,
$\limsup_{n\to\infty}\pi(\mathcal{E}(\delta_{n}))\leq\pi^{e}(F).$
(ii) Suppose $\liminf_{n\to\infty}\delta_{n}^{T(\mathcal{E}(\delta_{n}))}=1$. By passing to a convergent subsequence, we obtain
$\lim_{n\to\infty}\delta_{n}^{T(\mathcal{E}(\delta_{n}))}=\lim_{n\to\infty}q_{n}(\underline{\theta})=1,$
where $q_{n}(\underline{\theta})$ denotes the discounted allocation probability for type $\underline{\theta}$ in equilibrium $\mathcal{E}(\delta_{n})$. Since $q_{n}(\underline{\theta})\to 1$, it follows that $q_{n}(\theta)$ converges uniformly to $1$ for all $\theta$, because $q_{n}(\theta)\geq q_{n}(\underline{\theta})$. Moreover, in any equilibrium in which the lowest type is served, sequential rationality implies $x_{n}(\underline{\theta})\geq x^{e}(\underline{\theta})$. The incentive-compatibility constraint therefore implies
$q_{n}(\theta)x_{n}(\theta)\geq q_{n}(\underline{\theta})x^{e}(\underline{\theta}).$
As $n\to\infty$, the allocation converges to full participation, with $q=1$ and allocations bounded below by $x^{e}(\underline{\theta})$. Consequently, the seller’s payoff is asymptotically bounded above by the static benchmark $\pi^{e}(F)$:
$\limsup_{n\to\infty}\pi(\mathcal{E}(\delta_{n}))\leq\pi^{e}(F).$
The following corollary shows that the payoff upper bound $\pi(F)$ is generally unattainable.
Corollary 6.
Suppose Assumption 3 holds. If $\underline{x}<x^{e}(\underline{\theta})$ and the distribution $F$ has an atom at $\underline{\theta}$, then
$\limsup_{n\to\infty}\pi(\mathcal{E}(\delta_{n}))\;<\;\pi(F).$
Proof of Corollary 6.
Under these conditions, the revenue-maximizing static mechanism characterized in Lemma 3 prescribes an allocation $x^{*}(\underline{\theta})$ satisfying $x^{*}(\underline{\theta})<x^{e}(\underline{\theta})$, with trade probability $q(\underline{\theta})=1$.
Consider any sequence of equilibria $\{\mathcal{E}(\delta_{n})\}$. We analyze the outcome, possibly along a subsequence, for the lowest type $\underline{\theta}$:
  1. 1.
    Case 1: If the market does not clear in the limit, that is, if the probability of trade vanishes, then $\lim_{n\to\infty}q_{n}(\underline{\theta})<1$. Since the static optimum requires $q(\underline{\theta})=1$ and yields positive surplus from this type, the resulting dynamic payoff is strictly lower.
  2. 2.
    Case 2: If the market clears in finite time, an adaptation of Lemma 8 to the continuum-type setting, using Lemma 9, implies that the allocation satisfies
    $\liminf_{n\to\infty}x_{n}(\underline{\theta})\geq x^{e}(\underline{\theta})\quad\text{almost everywhere}.$
In the static benchmark, the contribution of type $\underline{\theta}$ to total profit is uniquely maximized at $x^{*}(\underline{\theta})$. Since $F$ has an atom at $\underline{\theta}$ and the dynamic equilibrium allocation $x_{n}(\underline{\theta})$ is bounded away from the static optimizer $x^{*}(\underline{\theta})$, the total expected revenue in the dynamic game is strictly less than $\pi(F)$ in the limit. ∎
B.2 Relaxing Assumption 3
We further assume that the price space is discretized. Specifically, the price must be drawn from the grid
$\{u_{\min},\,u_{\min}+\Delta,\,u_{\min}+2\Delta,\,\dots,\,\overline{p}\},$
for some sufficiently large $\overline{p}$.  [^14] One can allow that when $x=\overline{x}$, the price may be chosen from a continuum. This exception ensures that the discretization does not force immediate market clearing on the equilibrium path.
We show that, in the discrete action space model, Theorem 1 and Theorem 2 remain valid under Assumptions 1 and 2. Since Theorem 2 relies on the off-path threat constructed in Theorem 1, it suffices to establish that Theorem 1 holds in this setting.
Proposition 3.
Under Assumptions 1 and 2, for any $\epsilon>0$, there exists a threshold $\underline{\delta}\in(0,1)$ such that for all $\delta\geq\underline{\delta}$, there exists an equilibrium. In this equilibrium, the seller clears the market immediately at $t=0$ by offering $(\overline{x},u_{\text{min}})$, and obtains revenue equal to $u_{\text{min}}$.
Proof.
We first consider a candidate for on-path behavior. Suppose that, on the equilibrium path, allocation levels are fixed at $(\overline{x},u_{\min})$. In this case, every type attains full efficiency and satisfies
$u(\overline{x},\theta)=u(x^{e}(\theta),\theta).$
Now consider a deviation $(x_{t},p_{t})$. Since $x_{t}<\overline{x}$, if $p_{t}\geq u_{\min}+\Delta$, then for sufficiently large $\delta$, every buyer type prefers the anticipated equilibrium outcome. If $p_{t}=u_{\min}$, the market clears immediately. Hence, for sufficiently large $\delta$, the conjectured behavior is sustained by a direct application of the one-shot deviation principle. ∎
Proposition 4.
For every $\epsilon>0$, there exist $\Delta$ and a discount factor $\underline{\delta}$ such that for all $\delta\geq\underline{\delta}$,
$\pi^{e}(F)-\epsilon\in SE(\Delta,\delta),$
where $SE(\Delta,\delta)$ denotes the set of seller payoffs sustainable in equilibrium.
Using the equilibrium outcome characterized in Proposition 3 as the off-path punishment, we can sustain the menu described in Lemma 4. The proof is analogous to that of Theorem 2 and is therefore omitted. Since the action space is discretized, the continuum may not be attainable for arbitrary $\Delta>0$.
B.3 Menu Offers
In the baseline model, the seller is restricted to offering a single allocation–price pair in each period. If this constraint is relaxed to allow the seller to post arbitrary menus, the strategic environment changes.
Corollary 7.
The unique equilibrium outcome features immediate market clearing at $t=0$, in which the seller offers the optimal static menu characterized in Lemma 3 and obtains the unconstrained commitment payoff $\pi(F)$.
In this scenario, the static mechanism from Lemma 3 simultaneously maximizes revenue and clears the market. [^15] Since the seller can adjust prices by an arbitrarily small $\epsilon$ to strictly clear the market, the supremum is achieved in equilibrium; otherwise, profitable deviations would exist. This result corresponds to the intratemporal price discrimination effect identified by Wang (1998), Hahn (2006), Board and Pycia (2014), and Mensch (2017), and formalized by Nava and Schiraldi (2019). In these settings, the ability to screen types within a single period restores monopoly power. Consequently, the absence of intertemporal commitment is immaterial when the optimal market-clearing outcome coincides with the commitment payoff.
By contrast, the baseline model rules out intratemporal discrimination. In this case, the maximal market-clearing profit is $u_{\min}$, which is strictly below the static commitment benchmark. This difference explains why the equilibrium outcomes and underlying economic forces in the present model differ from those in the existing literature.
B.4 Indeterminacy of Delay
The argument is as follows. Choose $\epsilon>0$ such that $\pi^{e}(F)-\epsilon>u_{\min}$. Take $\delta$ sufficiently large so that $\pi^{e}(F)-\epsilon$ can be sustained in equilibrium as characterized in Theorem 2.
Consider such an equilibrium. The seller initially remains silent by making offers that the buyer never accepts; if the seller deviates, the players switch to the equilibrium in Theorem 1. If the seller remains silent for $k$ periods such that $\delta^{k}\geq u_{\min}/(\pi^{e}(F)-\epsilon)$, then after $k$ periods the players play the equilibrium in Theorem 2. Since
$\delta^{k}(\pi^{e}(F)-\epsilon)\geq u_{\min},$
this strategy profile constitutes an equilibrium. Because $\delta^{k}<1$, there is delay in allocation, and this delay does not vanish even as $\delta\to 1$.
B.5 Payoff Upper Bound without Free Disposability
To show that $\pi^{e}(F)$ cannot be achieved without free disposability in general, we establish the following proposition.
Proposition 5.
Suppose Assumptions 1, 2, and 3 hold. Consider an environment with three discrete types, ordered $\theta_{1}>\theta_{2}>\theta_{3}$, each with prior probability $q_{i}>0$ such that $\sum_{i=1}^{3}q_{i}=1$. The buyer’s utility is given by
$w(x,\theta)=v(x)+\theta x.$
As $\delta\to 1$, the seller’s equilibrium payoff is strictly bounded away from $\pi^{e}(F)$ in any equilibrium.
Proof of Proposition 5.
We proceed in several steps. First, we show that if an offer $(x_{t},p_{t})$ clears the last remaining mass of type $\theta_{2}$, then it must also clear the entire remaining mass of type $\theta_{1}$ (assuming $\theta_{1}>\theta_{2}$) at any history.
We argue by contradiction. Suppose instead that type $\theta_{2}$ clears while a positive mass of type $\theta_{1}$ remains. Consider the subsequent history in which type $\theta_{1}$ and possibly type $\theta_{3}$ remain active. If type $\theta_{3}$ were also cleared by the offer, the market would clear completely, which trivially satisfies the claim.
Invoking Proposition 1, we infer that continuation play is characterized by a sequence of offers with allocation $x^{e}(\theta_{1})$, followed by a final offer $x^{e}(\theta_{3})$, possibly with delay. In particular, in the period preceding the final clearing, a positive mass of type $\theta_{1}$ clears, satisfying the indifference condition between the allocations $x^{e}(\theta_{1})$ and $x^{e}(\theta_{3})$.
Suppose the final offer $x^{e}(\theta_{3})$ occurs after a delay of $s$ periods relative to $t$. The incentive constraints at period $t$ imply
$w(x_{t},\theta_{1})-p_{t}$ $\leq\delta^{s}\left[w(x^{e}(\theta_{3}),\theta_{1})-w(x^{e}(\theta_{3}),\theta_{3})\right],$
$w(x_{t},\theta_{2})-p_{t}$ $\geq\delta^{s}\left[w(x^{e}(\theta_{3}),\theta_{2})-w(x^{e}(\theta_{3}),\theta_{3})\right],$
$w(x_{t},\theta_{3})-p_{t}$ $\leq\delta^{s}\left[w(x^{e}(\theta_{3}),\theta_{3})-w(x^{e}(\theta_{3}),\theta_{3})\right].$
By the single-crossing property, the unique allocation satisfying these three inequalities is
$x_{t}=\delta^{s}x^{e}(\theta_{3}),$
and all three inequalities bind. This implies that type $\theta_{3}$ is indifferent between purchasing and waiting at time $t$. By the tie-breaking assumption, the market therefore clears at time $t$, leading a contradiction. [^16] Even if type $\theta_{3}$ does not break indifference in favor of the seller, provided $\underline{x}\geq x^{e}(\theta_{3})$, the conclusion continues to hold because in that case the three inequalities cannot be satisfied simultaneously.
Define
$p^{*}=w(x^{e}(\theta_{2}),\theta_{2})-\delta\Big[w(x^{e}(\theta_{3}),\theta_{2})-w(x^{e}(\theta_{3}),\theta_{3})\Big].$
This price renders type $\theta_{2}$ indifferent between purchasing immediately and waiting to mimic type $\theta_{3}$ in the subsequent period, when the market clears.
Consider the set $S$ of all pairs $(x,p)$ that type $\theta_{2}$ prefers to accept while type $\theta_{3}$ prefers to wait in all equilibria. Define
$s^{*}=\sup_{(x,p)\in S}\big(w(x,\theta_{2})-p\big)$
to be the maximum possible surplus for $\theta_{2}$ among all these pairs. We claim that $s^{*}\leq w(x^{e}(\theta_{2}),\theta_{2})-p^{*}$. We argue by contradiction. Recall that if type $\theta_{2}$ clears completely, then type $\theta_{1}$ must also clear. Consequently, the equilibrium path in the subsequent period is uniquely determined. The seller must clear type $\theta_{3}$ with the offer $(x^{e}(\theta_{3}),w(x^{e}(\theta_{3}),\theta_{3}))$.
Now suppose that $s^{*}>w(x^{e}(\theta_{2}),\theta_{2})-p^{*}$. Then there exist $\epsilon>0$ and an offer $(x,p)$ such that the surplus satisfies
$w(x,\theta_{2})-p\in(s^{*}-\epsilon,s^{*}].$
If the discount factor and the gap satisfy
$s^{*}-\epsilon>\delta s^{*}\quad\text{and}\quad s^{*}-\epsilon>w(x^{e}(\theta_{2}),\theta_{2})-p^{*},$
the seller can construct a profitable deviation. By offering a slightly higher price, which reduces the surplus while keeping it above the continuation value, type $\theta_{2}$ and therefore type $\theta_{1}$ still clear immediately, followed by the clearing of type $\theta_{3}$. This deviation yields a strictly higher profit, contradicting optimality of the original strategy. Hence, it must be that $s^{*}\leq w(x^{e}(\theta_{2}),\theta_{2})-p^{*}$.
Observe that the offer $(x^{e}(\theta_{2}),p^{*})$ yields a surplus of exactly $w(x^{e}(\theta_{2}),\theta_{2})-p^{*}$ to type $\theta_{2}$. We argue that, regardless of how small the initial mass of type $\theta_{2}$ is, this offer must appear on the equilibrium path when $\delta$ is sufficiently close to $1$, which makes screening type $\theta_{2}$ profitable.
Consider the equilibrium dynamics. There must exist a period in which type $\theta_{2}$ is fully cleared; denote this period by $t$. There are four possible scenarios.
  1. 1.
    Clear $\theta_{2}$, then clear $\theta_{3}$. The seller clears type $\theta_{2}$ in the penultimate period, followed by type $\theta_{3}$ in the final period. This structure implies that the offer $(x^{e}(\theta_{2}),p^{*})$ appears on the equilibrium path. Otherwise, $(x^{e}(\theta_{2}),p^{*})$ would constitute a profitable deviation.
  2. 2.
    Clear $\theta_{1}$ and $\theta_{2}$, then clear $\theta_{3}$. The seller clears types $\theta_{1}$ and $\theta_{2}$ simultaneously in the penultimate period, followed by type $\theta_{3}$ in the final period. This structure again implies that the offer $(x^{e}(\theta_{2}),p^{*})$ appears on the equilibrium path. Otherwise, $(x^{e}(\theta_{2}),p^{*})$ would constitute a profitable deviation, for the same reason as above.
  3. 3.
    Pool $\theta_{2}$ and $\theta_{3}$. Suppose the seller leaves a small mass of $\theta_{2}$ that is pooled with $\theta_{3}$ and cleared in one single period. The remaining mass of $\theta_{2}$ is sufficiently small that the seller prefers to clear the market immediately, including type $\theta_{3}$, rather than delay sales to $\theta_{3}$ when facing only types $\theta_{2}$ and $\theta_{3}$.
    This implies that there exists a period $s$ prior to this final pooling period in which some mass of type $\theta_{2}$ is cleared, and period $s$ is the penultimate period in which $\theta_{2}$ is served. At period $s$, type $\theta_{1}$ must also be cleared. Otherwise, the same reasoning as in the previous cases applies: the final period would clear $\theta_{2}$ and $\theta_{3}$ together rather than only $\theta_{3}$, yet the relevant inequalities would be unchanged.
    By the same argument as above, no mass of $\theta_{1}$ can remain at period $s$. Therefore, the offer at period $s$ must be $(x^{e}(\theta_{2}),p^{*})$; otherwise, there exists a profitable deviation.
  4. 4.
    Clear $\theta_{1},\theta_{2},\theta_{3}$ simultaneously. In this case, the remaining mass of $\theta_{2}$ must be sufficiently small. Otherwise, the seller would prefer to separate $\theta_{1}$, $\theta_{2}$, and $\theta_{3}$ rather than clear them jointly, for the same reason as before. This implies that there exists a period $s$ prior to the final polling period in which some mass of type $\theta_{2}$ is cleared. At period $s$, type $\theta_{1}$ must also be cleared. Otherwise, the same reasoning as in the previous cases applies, which leads to a contradiction.
We have established that the offer $(x^{e}(\theta_{2}),p^{*})$ must appear on the equilibrium path. To attain the seller payoff characterized in Lemma 4, there must be asymptotically no delay, since the optimal static menu requires the allocation probability for the lowest type to satisfy $q(\theta_{3})=1$.
Recall that the static mechanism in Lemma 4 distorts the allocation for type $\theta_{2}$ to $x^{*}(\theta_{2})$, with a corresponding price, in order to minimize the information rent of type $\theta_{1}$. In the dynamic game, because type $\theta_{3}$ is not asymptotically delayed, no type experiences delay. Consequently, type $\theta_{1}$ can mimic type $\theta_{2}$ and obtain the offer $(x^{e}(\theta_{2}),p^{*})$. Since $x^{e}(\theta_{2})>x^{*}(\theta_{2})$, this deviation yields a strictly larger information rent for type $\theta_{1}$ than under the static mechanism. Therefore, the rents extracted from type $\theta_{1}$, net of the rents from type $\theta_{2}$, are strictly smaller than in the static benchmark. This proves the proposition.
Appendix C Institutional Evidence
C.1 Importance of B2B Digital Goods Market
The market for Business-to-Business (B2B) digital goods—including enterprise software, financial data, and AI models—is a central component of the modern economy. Market leaders in this sector exhibit high profitability (Guellec and Paunov, 2017). Direct evidence comes from the disclosures of publicly traded corporations. Specifically, annual Form 10-K filings provide an audited decomposition of revenue streams.
An examination of filings from industry leaders shows a consistent pattern: the enterprise sector accounts for a large share of profits. Microsoft Corporation’s enterprise-focused segments constitute the primary drivers of its profitability. The Intelligent Cloud segment generated \$87.9 billion in revenue in fiscal year 2023. The Productivity and Business Processes segment derives most of its revenue from commercial clients, with Office Commercial products and cloud services contributing \$48.7 billion [^17] See Microsoft Corporation (2023).. Salesforce reported \$34.9 billion in revenue for fiscal year 2024, generated from B2B subscriptions to its cloud-based software [^18] See Salesforce, Inc. (2024).. Adobe Inc.’s Digital Media segment generated \$14.2 billion in 2023, a substantial portion of which came from enterprise subscriptions. The Digital Experience segment is an exclusively B2B division providing marketing, analytics, and e-commerce services, generating \$4.9 billion in revenue [^19] See Adobe Inc. (2023)..
This underscores the importance of understanding the market dynamics in this industry.
C.2 Single-offer Protocol
The model assumes that the seller makes a single take-it-or-leave-it offer in each period. Although the confidential nature of B2B negotiations often precludes direct observation of the bargaining process, we discuss the Amazon Web Services “Private Offer” feature as an institutional commitment device.
The institutional design of major B2B marketplaces, including the AWS Marketplace, is consistent the single-offer protocol. On these platforms, Independent Software Vendors (ISVs) such as Snowflake and data providers such as Reuters transact with enterprise buyers. To facilitate these transactions, the marketplace employs a “Private Offer” mechanism (Amazon Web Services (2025)).
This mechanism serves as an institutional commitment device. The seller issues a formal, binding proposal to a specific buyer that is valid for a fixed duration. If the buyer rejects the offer, the seller may issue a subsequent proposal. Crucially, each offer consists of a single, indivisible set of terms, including the software tier, user count, and price. The buyer therefore faces a binary acceptance decision over a bundled proposal. This design is consistent with the model’s structural assumption: regardless of informal pre-negotiation, the platform’s architecture requires the seller to commit to one binding offer at a time. The fact that the industry-standard instrument for executing customized transactions takes the form of a single offer is consistent with for the framework.